### v1

# An Effective Deligne’s equidistribution theorem

Abstract.
We prove an Erdős–Turán type inequality for compact Lie groups, from which we deduce an effective version of Deligne’s equidistribution theorem.
Key words and phrases:
Deligne’s equidistribution theorem, Erdős–Turán inequality, Weyl integration formula, Weyl character formula.
2020 Mathematics Subject Classification:
11K38, 22E46, 14F20
1. Introduction
Let $\mathbf{F}_{q}$ be a finite field of characteristic $p$ with $q$ elements, $X$ a smooth geometrically connected algebraic curve over $\mathbf{F}_{q}$, $\eta$ the generic point of $X$, $\ell$ a prime number distinct from $p$, and $\mathcal{F}$ a lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$. We have an exact sequence of fundamental groups
$1\to\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})\to\pi_{1}(X,\bar{\eta})\to\mathrm{Gal}(\overline{\mathbf{F}}_{q}/\mathbf{F}_{q})\to 1$
The lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf $\mathcal{F}$ on $X$ defines a representation
$\rho_{\mathcal{F}}:\pi_{1}(X,\bar{\eta})\to\mathrm{GL}(\mathcal{F}_{\bar{\eta}}).$
The geometric monodromy group (resp. arithmetic monodromy group) of $\mathcal{F}$ is defined to be the Zariski closure of the image of $\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})$ (resp. $\pi_{1}(X,\bar{\eta}))$ in $\mathrm{GL}(\mathcal{F}_{\bar{\eta}})$.
Fix an isomorphism $\iota:\overline{\mathbf{Q}}_{\ell}\stackrel{{\scriptstyle\cong}}{{\to}}\mathbf{C}$. Throughout this paper, we assume $\mathcal{F}$ is punctually $\iota$-pure of weight 00, that is, for any closed point $x$ in $X$ and any eigenvalue $\lambda$ of the action of the geometric Frobenius element $F_{x}$ on $\mathcal{F}_{\bar{x}}$, we have $|\iota(\lambda)|=1$. Then by theorems of Deligne and Grothendieck [De80, 3.4.1(iii), 1.3.9], the identity component $G^{0}$ of the geometric monodromy group $G$ is semisimple. Via the isomorphism $\iota$, the algebraic group $G$ defines a complex semisimple Lie group $G_{\mathbf{C}}$. The category of finite dimensional $\overline{\mathbf{Q}}_{\ell}$-representations of $G$ is equivalent to that of finite dimensional complex representations of $G_{\mathbf{C}}$.
We assume $G$ is connected throughout this paper. Let $\mathfrak{g}$ be the Lie algebra of $G_{\mathbf{C}}$. Fix a Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{g}$, and let $R$ be the root system. Choose an ordering $R=R^{+}\cup R^{-}$ on $R$. For each ${\boldsymbol{\alpha}}\in R$, let $H_{\boldsymbol{\alpha}}\in\mathfrak{h}$ be the element in $[\mathfrak{g}_{\boldsymbol{\alpha}},\mathfrak{g}_{-{\boldsymbol{\alpha}}}]$ satisfying ${\boldsymbol{\alpha}}(H_{\boldsymbol{\alpha}})=2$, and let $\mathfrak{h}_{\mathbf{R}}$ be the real subspace of $\mathfrak{h}$ spanned by $H_{\boldsymbol{\alpha}}$ $(\boldsymbol{\alpha}\in R)$. There exists a real subalgebra $\mathfrak{g}_{0}$ of $\mathfrak{g}$ such that
$\mathfrak{g}_{0}=i\mathfrak{h}_{\mathbf{R}}\bigoplus\bigoplus_{{\boldsymbol{\alpha}}\in R^{+}}\Big{(}\mathfrak{g}_{0}\cap(\mathfrak{g}_{\boldsymbol{\alpha}}\oplus\mathfrak{g}_{-{\boldsymbol{\alpha}}})\Big{)},$
that $\mathfrak{g}_{0}$ is the Lie algebra of a maximal compact subgroup ${G_{\mathbf{R}}}$ of $G$, and that $i\mathfrak{h}_{\mathbf{R}}$ is the Lie algebra of a maximal torus $T$ of ${G_{\mathbf{R}}}$. The category of finite dimensional complex representations of $G_{\mathbf{C}}$ is equivalent to that of finite dimensional complex representations of $G_{\mathbf{R}}$.
For simplicity, we assume $\rho_{\mathcal{F}}(\pi_{1}(X,\bar{\eta}))\subset G$ as in [Ka88, 3.3], that is, the arithmetic monodromy group coincides with the geometric monodromy group. Then for every closed point $x$ in $X$, the geometric Frobenius conjugacy class $F_{x}$ in $\pi_{1}(U,\bar{\eta})^{\natural}$ defines an element in $G^{\natural}$ and an element in $G^{\natural}_{\mathbf{C}}$, where the set of conjugacy classes of a group $H$ is denoted by $H^{\natural}$. Since we assume $\mathcal{F}$ is punctually $\iota$-pure of weight 00, the semisimple part of $\rho_{\mathcal{F}}(F_{x})$ defines an element in $G_{\mathbf{R}}^{\natural}$ which we denote by $\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}$. Provide $G_{\mathbf{R}}^{\natural}$ with the quotient topology induced from $G_{\mathbf{R}}$, and let $\mu_{G_{\mathbf{R}}^{\natural}}$ be the measure on $G_{\mathbf{R}}^{\natural}$ induced by the Haar measure $\mu_{G_{\mathbf{R}}}$ on $G_{\mathbf{R}}$. For any integrable function $f$ on $G_{\mathbf{R}}^{\natural}$, we have
$\int_{G_{\mathbf{R}}^{\natural}}f\,\mathrm{d}\mu_{G_{\mathbf{R}}^{\natural}}=\int_{G_{\mathbf{R}}}f\,\mathrm{d}\mu_{G_{\mathbf{R}}}.$
Deligne proves the following theorem ([De80, 3.5.3], [Ka88, 3.6]).
Theorem 1.1 (Deligne).
Let $\mathcal{F}$ be a lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$ punctually $\iota$-pure of weight 00, and let $G$ be its geometric monodromy group. Suppose the arithmetic monodromy group of $\mathcal{F}$ coincides with $G$. For any positive integer $m$, let $X(\mathbf{F}_{q^{m}})$ be the set of $\mathbf{F}_{q^{m}}$-points in $X$. For any domain $D$ in $G_{\mathbf{R}}^{\natural}$ so that the boundary of $D$ has measure 00, we have
$\lim_{m\to\infty}\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D\}|}{|X(\mathbf{F}_{q^{m}})|}=\mu_{G_{\mathbf{R}}^{\natural}}(D).$
The main result of this paper is an effective version of the above theorem.
Theorem 1.2.
Let $\mathcal{F}$ be a lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$ punctually $\iota$-pure of weight 00, and let $G$ be its geometric monodromy group. Suppose $G$ is connected and the arithmetic monodromy group of $\mathcal{F}$ coincides with $G$. Then for any small box $D$ in $G^{\natural}_{\mathbf{R}}$, we have
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D\}|}{|X(\mathbf{F}_{q^{m}})|}=\mu_{G_{\mathbf{R}}^{\natural}}(D)+O\Big{(}q^{-\frac{m}{2(|R^{+}|+n)}}\Big{)},$
where $n=\mathrm{dim}_{\mathbf{C}}\mathfrak{h}$ is the rank of the Lie group, and the constant implied by $O$ depends only on $X$ and $\mathcal{F}$.
We will describe small boxes in Definition 1.3 below. We deduce the above theorem from an Erdős-Turán type inequality for compact Lie groups, which is of independent interest and we now describe. Let
$N(T):=\{g\in{G_{\mathbf{R}}}:gTg^{-1}=T\}$
be the normalizer of the maximal torus $T$ in ${G_{\mathbf{R}}}$, and let $W:=N(T)/T$ be the Weyl group. It acts on $T$ by
$W\times T\to T,\quad(gT,x)\mapsto gxg^{-1}.$
Hence $W$ acts on the Lie algebra $\mathfrak{h}_{\mathbf{R}}$ and its dual $\mathfrak{h}_{\mathbf{R}}^{*}$. Provide $\mathfrak{h}_{\mathbf{R}}$ and $\mathfrak{h}_{\mathbf{R}}^{*}$ with the inner product induced by the Killing form. The Weyl group can be identified with the subgroup of $\mathrm{GL}(\mathfrak{h}_{\mathbf{R}})$ generated by reflections with respect to $H_{\boldsymbol{\alpha}}^{\perp}$ $({\boldsymbol{\alpha}}\in R)$, and with the subgroup of $\mathrm{GL}(\mathfrak{h}_{\mathbf{R}}^{*})$ generated by reflections with respect to ${\boldsymbol{\alpha}}^{\perp}$ $({\boldsymbol{\alpha}}\in R)$. Let $T/W$ be the quotient space. The map $T\to G_{\mathbf{R}}^{\natural}$ sending each element in $T$ to its conjugacy class in $G_{\mathbf{R}}$ defines a homeomorphism
$\kappa:T/W\stackrel{{\scriptstyle\cong}}{{\to}}G_{\mathbf{R}}^{\natural}$
by [BtD85, IV Proposition 2.6] if we provide $T/W$ and $G_{\mathbf{R}}^{\natural}$ with the quotient topology induced from $T$ and $G_{\mathbf{R}}$, respectively.
Any representation $G\to\mathrm{GL}(V)$ gives rise to a representation of $\mathfrak{g}\to\mathfrak{gl}(V)$ such that the following diagram commutes:
$\begin{array}[]{ccccc}2\pi i\mathfrak{h}_{\mathbf{R}}&\hookrightarrow&\mathfrak{g}_{0}&\to&\mathfrak{gl}(V)\\ {\scriptstyle\exp}\downarrow&&{\scriptstyle\exp}\downarrow&&{\scriptstyle\exp}\downarrow\\ T&\hookrightarrow&G_{0}&\to&\mathrm{GL}(V).\end{array}$
Let ${\boldsymbol{\lambda}}\in\mathfrak{h}^{*}$ be a weight of the representation. Then there exists a nonzero vector $v\in V$ such that $Hv={\boldsymbol{\lambda}}(H)v$ for all $H\in\mathfrak{h}$. We thus have
$\exp(2\pi iH)(v)=e^{2\pi i{\boldsymbol{\lambda}}(H)}v.$
The map $\exp:2\pi i\mathfrak{h}_{\mathbf{R}}\to T$ is an epimormphism. Let $2\pi i\Gamma$ be its kernel. We call $\Gamma$ the integral lattice. If $H\in\Gamma$, then we have $e^{2\pi i{\boldsymbol{\lambda}}(H)}v=v$, and hence $e^{2\pi i{\boldsymbol{\lambda}}(H)}=1$. So we have
${\boldsymbol{\lambda}}(\Gamma)\subset\mathbf{Z}.$
Define the lattice of integral forms $I^{*}$ to be
$I^{*}:=\{{\boldsymbol{\lambda}}\in\mathfrak{h}_{\mathbf{R}}^{*}:\,{\boldsymbol{\lambda}}(\Gamma)\subset\mathbf{Z}\}.$
For any ${\boldsymbol{\lambda}}\in I^{*}$, define a character
$e({\boldsymbol{\lambda}}):T\to S^{1}:=\{z\in\mathbf{C}:\,|z|=1\}$
which maps $\exp(2\pi iH)\in T$ to $e^{2\pi i{\boldsymbol{\lambda}}(H)}\in S^{1}$ for every $H\in\mathfrak{h}_{\mathbf{R}}$. We thus get an isomorphism
$I^{*}\stackrel{{\scriptstyle\cong}}{{\to}}\mathrm{Hom}(T,S^{1}),\quad\boldsymbol{\lambda}\mapsto e(\boldsymbol{\lambda}).$
In particular, for any root ${\boldsymbol{\alpha}}\in R$, we have a character $e({\boldsymbol{\alpha}})$ of $T$. Define the weight lattice to be
$\Lambda:=\{{\boldsymbol{\lambda}}\in\mathfrak{h}_{\mathbf{R}}^{*}:\,{\boldsymbol{\lambda}}(H_{\boldsymbol{\alpha}})\in\mathbf{Z}\hbox{ for all }{\boldsymbol{\alpha}}\in R\}.$
We have
$\sum_{\boldsymbol{\alpha}\in R}\mathbf{Z}H_{\boldsymbol{\alpha}}~{}\subseteq$ $~{}\Gamma~{}\subseteq~{}\{H\in\mathfrak{h}_{\mathbf{R}}:\,\boldsymbol{\alpha}(H)\in\mathbf{Z}\hbox{ for all }\boldsymbol{\alpha}\in R\},$
$\Lambda~{}\supseteq$ $~{}I^{*}~{}\supseteq~{}\sum_{\boldsymbol{\alpha}\in R}\mathbf{Z}{\boldsymbol{\alpha}}.$
Definition 1.3.
Fix a basis $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}$ for the lattice $\Gamma$. A subset of $T$ of the form
$E=\Big{\{}\exp\Big{(}2\pi i\sum_{j=1}^{n}t_{j}\mathbf{e}_{j}\Big{)}:\,t_{j}\in I_{j}\Big{\}}$
for some intervals $I_{j}$ of lengths $<1$ is called a box in $T$. We say the box $E$ is small if for any nonidentity element $\sigma$ in the Weyl group $W$, $E\cap\sigma(E)$ is empty. A subset $D$ of $T/W$ (resp. $G_{\mathbf{R}}^{\natural}$) is called a small box if it is the image of a small box $E$ in $T$ under the map
$T\to T/W\quad(\mathrm{resp.}~{}T\to T/W\underset{\cong}{\stackrel{{\scriptstyle\kappa}}{{\to}}}G_{\mathbf{R}}^{\natural}).$
Remark 1.4.
With the above notations, since the lengths of the intervals $I_{j}$ are $<1$, the map $\exp(2\pi i\hbox{-}):\mathfrak{h}_{R}\to T$ identifies the set $I_{1}\times\cdots\times I_{n}$ with its image $E$ in $T$. Suppose furthermore that $E\cap\sigma(E)$ is empty for every nonidentity element $\sigma\in W$. Then the projection $T\to T/W$ identifies $E$ with its image $D$ in $T/W$, and the inverse image of $D$ in $T$ is $\bigsqcup_{\sigma\in W}\sigma(E)$.
Let
:=$\mathfrak{C}:=\{{\boldsymbol{\lambda}}\in\mathfrak{h}_{\mathbf{R}}^{*}:\;{\boldsymbol{\lambda}}(H_{\boldsymbol{\alpha}})\in\mathbf{R}_{\geq 0}\hbox{ forall }{\boldsymbol{\alpha}}\in R^{+}\}$
be the dominant Weyl chamber in $\mathfrak{h}_{R}^{*}$. For any ${\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C}$, define
$A_{\boldsymbol{\lambda}}:=\sum_{\sigma\in W}\mathrm{sgn}(\sigma)e(\sigma({\boldsymbol{\lambda}})),$
where for any $\sigma\in W$, $\mathrm{sgn}(\sigma)$ is defined to be determinant of $\sigma$ considered as a linear operator on $\mathfrak{h}_{\mathbf{R}}^{*}$. Its values are ±$\pm 1$ since $\sigma$ is a composite of reflections. Let
${\boldsymbol{\rho}}=\frac{1}{2}\sum_{{\boldsymbol{\alpha}}\in R^{+}}{\boldsymbol{\alpha}}$
be half of the sum of positive roots. Then $\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}$ is a trigonometric polynomial function on $T$ invariant under the action of $W$ ([BtD85, VI 1.6]). It defines a continuous function on $T/W\cong G^{\natural}_{\mathbf{R}}$ which we also denote by $\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}$.
Let $\boldsymbol{\boldsymbol{\alpha}}_{1},\ldots,\boldsymbol{\boldsymbol{\alpha}}_{n}\in R^{+}$ be all the simple roots. The fundamental weights $\boldsymbol{\omega}_{1},\ldots,\boldsymbol{\omega}_{n}$ in the weight lattice $\Lambda$ are defined by
$\boldsymbol{\omega}_{i}(H_{\boldsymbol{\boldsymbol{\alpha}}_{j}})=\delta_{ij}.$
Any $\boldsymbol{\boldsymbol{\lambda}}\in\mathfrak{h}^{*}_{\mathbf{R}}$ can be written as $\boldsymbol{\boldsymbol{\lambda}}=\sum_{j=1}^{n}{\lambda}_{j}\boldsymbol{\omega}_{j}$ with $\lambda_{j}\in\mathbf{R}$, and $\boldsymbol{\boldsymbol{\lambda}}$ lies in $\Lambda\cap\mathfrak{C}$ if and only if ${\lambda}_{j}$ are nonnegative integers. The following is a generation of the Erdős–Turán inequality ([Mo94, Chapter 1, Corollary 1.1]) to compact Lie groups.
Theorem 1.5.
For any sequence $x_{1},\ldots,x_{N}\in G_{\mathbf{R}}^{\natural}$, any small box $D$ in $G_{\mathbf{R}}^{\natural}$ and any positive integer $M$, we have
|$\frac{|\{1\leq i\leq N:x_{i}\in D\}|}{N}-\mu_{G_{\mathbf{R}}^{\natural}}(D)$
$\ll$ $\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{0\}\\ 0\leq\lambda_{j}\leq M\end{subarray}}\frac{1}{\|\boldsymbol{\lambda}\|+1}\Big{|}\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})\Big{|}+\frac{1}{M},$
where $\|\boldsymbol{\lambda}\|=\sqrt{\sum_{i=1}^{n}\lambda_{i}^{2}},$ and the constant implied by $\ll$ depends only on $G$.
Remark 1.6.
For $G=\mathrm{SL}_{2}$, Theorem 1.5 is due to Niederreiter [Ni91, Lemma 3]. Rouse and Thorner [RT17, Lemma 3.1] give another proof of Niederreiter’s theorem using the Beurling and Selberg majorizing and minorizing functions. We prove Theorem 1.5 using Cochrane’s $n$-dimensional version of the Beurling and Selberg functions ([Co85]).
Let $\mathcal{F}^{\prime}$ be another lisse $\overline{\mathbf{Q}}_{\ell}$-sheaves on $X$ punctually $\iota$-pure of weight 00, $\rho_{\mathcal{F}^{\prime}}:\pi_{1}(X,\bar{\eta})\to\mathrm{GL}(\mathcal{F}^{\prime}_{\bar{\eta}})$ the corresponding representation, $G^{\prime}$ the geometric monodromy group, $G^{\prime}_{\mathbf{R}}$ a maximal compact subgroup of $G^{\prime}_{\mathbf{C}}$, $\mu_{G^{\prime\natural}_{\mathbf{R}}}$ the measure on $G^{\prime\natural}_{\mathbf{R}}$ induced by the Haar measure on $G^{\prime}_{\mathbf{R}}$. We have the following theorem on joint distribution.
Theorem 1.7.
Let $\mathcal{F}$ and $\mathcal{F}^{\prime}$ be two lisse $\overline{\mathbf{Q}}_{\ell}$-sheaves punctually $\iota$-pure of weight 00. Suppose their geometric mondromy groups $G$ and $G^{\prime}$ are connected and coincide with their arithmetic monodromy groups, respectively. Suppose furthermore that for any nontrivial irreducible $\overline{\mathbf{Q}}_{\ell}$-representations $\Gamma$ of $G$ and $\Gamma^{\prime}$ of $G^{\prime}$, the representation $(\Gamma\circ\rho_{\mathcal{F}})\otimes(\Gamma^{\prime}\circ\rho_{\mathcal{F}^{\prime}})$ has no nonzero $\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})$-invariant. Then for any small boxes $D$ in $G^{\natural}_{\mathbf{R}}$ and $D^{\prime}$ in $G^{\prime\natural}_{\mathbf{R}}$, we have
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\,\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D,\,\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{s}s}\in D^{\prime}\}|}{|X(\mathbf{F}_{q^{m}})|}$
$=$ $\mu_{G^{\natural}_{\mathbf{R}}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})+O\Big{(}q^{-\frac{m}{2(|R^{+}|+|R^{\prime+}|+n+n^{\prime}-1)}}\Big{)},$
where $n$ and $n^{\prime}$ are the dimensions of the Cartan subalgebras of $\mathfrak{g}$ and $\mathfrak{g}^{\prime}$, and the constant implied by $O$ depends only on $X$, $\mathcal{F}$ and $\mathcal{F}^{\prime}$.
The paper is organized as follows. In Section 2, we prove Theorem 1.5 using the Weyl integration formula, the Weyl character formula, and Cochrane’s theorem. In Section 3, we deduce Theorems 1.2 from Theorem 1.5 and Deligne’s theorem (the Weil conjecture). In Section 4, we study joint distribution and prove Theorem 1.7. As an application, we prove a joint distribution theorem for the Kloosterman sums and the Airy sums.
Acknowledgements
We would like to thank Winnie Li and Peng Shan for inspiring discussions. LF is supported by 2021YFA1000700, 2023YFA1009703 and NSFC12171261. YKL is supported by GRF (No. 17303619, 17307720). PX is supported by NSFC (No. 12025106).
2. Proof of Theorem 1.5
Let $\mu_{{G_{\mathbf{R}}}}$ and $\mu_{T}$ be the Haar measures on ${G_{\mathbf{R}}}$ and $T$, respectively. Recall that $\mu_{G_{\mathbf{R}}^{\natural}}$ is the measure on $G_{\mathbf{R}}^{\natural}$ induced by $\mu_{G_{\mathbf{R}}}$. Let $\mu_{T/W}$ be the measure on $T/W$ induced by $\mu_{T}$. Elements in the Weyl group $W$ permute $\{e({\boldsymbol{\alpha}})\}_{{\boldsymbol{\alpha}}\in R}$. So $\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))$ defines a function on $T/W$.
Proposition 2.1.
Via the homeomorphism $\kappa:T/W\stackrel{{\scriptstyle\cong}}{{\to}}G_{\mathbf{R}}^{\natural}$, the measure $\mu_{G_{\mathbf{R}}^{\natural}}$ is identified with the measure $\big{(}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\big{)}\mu_{T/W}$ on $T/W$.
Proof.
Let $\pi:G_{\mathbf{R}}\to G_{\mathbf{R}}^{\natural}$ be the canonical map. For any integrable function $f$ on $G_{\mathbf{R}}^{\natural}$, by the Weyl integration formula [BtD85, IV 1.11, VI 1.8], we have
$\int_{G_{\mathbf{R}}^{\natural}}f\mathrm{d}\mu_{G_{\mathbf{R}}^{\natural}}$ $=\int_{G_{\mathbf{R}}}f\circ\pi\,\mathrm{d}\mu_{G_{\mathbf{R}}}=\frac{1}{|W|}\int_{T}(f\circ\pi|_{T})\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}$
$=\frac{1}{|W|}\int_{T/W}(f\circ\kappa)\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T/W}.$
Proposition 2.2.
The set $\{\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}:{\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C}\}$ is an orthonormal basis for the Hilbert space $L^{2}\big{(}T/W,\big{(}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\big{)}\mu_{T/W}\big{)}.$
Proof.
By [BtD85, Theorem VI 1.7 (ii)], any irreducible representation of $G_{\mathbf{R}}$ has its highest weight lying in $I^{*}\cap\mathfrak{C}$, and for any ${\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C}$, there exists an irreducible representation $\Gamma_{\boldsymbol{\lambda}}$ of $G_{\mathbf{R}}$ unique up to isomorphism with the highest weight ${\boldsymbol{\lambda}}$. By the Peter-Weyl theorem ([BtD85, Theorem III 3.1]), the characters of $\Gamma_{\boldsymbol{\lambda}}$ $({\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C})$ form an orthonormal basis of $L^{2}(G_{\mathbf{R}}^{\natural},\mu_{G_{\mathbf{R}}^{\natural}})$. Identify $(G_{\mathbf{R}}^{\natural},\mu_{G_{\mathbf{R}}^{\natural}})$ with $\Big{(}T/W,\big{(}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\big{)}\mu_{T/W}\Big{)}$. By the Weyl character formula [BtD85, VI 1.7 (ii)-(iii)], the function on $T/W$ defined by the character of $\Gamma_{\boldsymbol{\lambda}}$ is exactly $\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}.$
As in the introduction, we write $\boldsymbol{\lambda}\in\Lambda$ as
$\boldsymbol{\lambda}=\sum_{j=1}^{n}\lambda_{j}\boldsymbol{\omega}_{j},$
where $\boldsymbol{\omega}_{1},\ldots,\boldsymbol{\omega}_{n}$ are the fundamental weights. Define
$\|\boldsymbol{\lambda}\|:=\sqrt{\lambda_{1}^{2}+\cdots+\lambda_{n}^{2}},\quad\mathbf{N}(\boldsymbol{\lambda}):=\|\boldsymbol{\lambda}\|+1.$
Lemma 2.3.
Let $\sigma$ be an element in the Weyl group $W$, and let $\boldsymbol{\mu}\in\Lambda$. There exists a constant $C>0$ such that for any $\boldsymbol{\lambda}\in\Lambda$, we have
$\mathbf{N}(\sigma(\boldsymbol{\lambda})+\boldsymbol{\mu})\geq C\mathbf{N}(\boldsymbol{\lambda})$
Proof.
Let $\boldsymbol{\nu}=\sigma(\boldsymbol{\lambda})+\boldsymbol{\mu}$. We have $\boldsymbol{\lambda}=\sigma^{-1}(\boldsymbol{\nu}-\boldsymbol{\mu}).$ Let
$\|\sigma^{-1}\|=\sup_{\|\boldsymbol{\alpha}\|=1}\|\sigma^{-1}(\boldsymbol{\alpha})\|.$
We then have
$\mathbf{N}(\boldsymbol{\lambda})=$ $\|\boldsymbol{\lambda}\|+1=\|\sigma^{-1}(\boldsymbol{\nu}-\boldsymbol{\mu})\|+1\leq\|\sigma^{-1}\|\|\boldsymbol{\nu}\|+\|\sigma^{-1}(\boldsymbol{\mu})\|+1$
$\leq$ $\max\Big{(}\|\sigma^{-1}\|,\|\sigma^{-1}(\boldsymbol{\mu})\|+1\Big{)}\mathbf{N}(\boldsymbol{\nu}).$
Our assertion follows. ∎
Lemma 2.4.
Let $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}$ be the basis of the integral lattice $\Gamma$ in Definition 1.3. It is also a basis for $\mathfrak{h}_{\mathbf{R}}$. Let $\{\mathbf{e}^{*}_{1},\ldots,\mathbf{e}^{*}_{n}\}$ be the dual basis for $\mathfrak{h}^{*}_{\mathbf{R}}$. For any $\boldsymbol{\lambda}\in\mathfrak{h}^{*}_{\mathbf{R}}$, write
$\boldsymbol{\lambda}=\sum_{j=1}^{n}\check{\lambda}_{j}\mathbf{e}^{*}_{j}.$
There exists a constant $c>0$ such that for any positive integer $L$ and any $\sigma\in W$, if $\max_{j}|\check{\lambda}_{j}|\leq L$, then $\max_{j}|(\sigma(\boldsymbol{\lambda}))_{j}|\leq cL$, where $\sigma(\boldsymbol{\lambda})=\sum_{j=1}^{n}(\sigma(\boldsymbol{\lambda}))_{j}\boldsymbol{\omega}_{j}.$
Proof.
Let $A({\sigma})=(a_{ij}(\sigma))$ be the matrix defined by $\sigma(\mathbf{e}_{j}^{*})=\sum_{i=1}^{n}\boldsymbol{\omega}_{i}a_{ij}(\sigma)$. We then have
$(\sigma(\boldsymbol{\lambda}))_{i}=\sum_{j=1}^{n}a_{ij}(\sigma)\check{\lambda}_{j}.$
It suffices to take $c=\max_{i,\sigma}\sum_{j=1}^{n}|a_{ij}(\sigma)|$. ∎
Lemma 2.5.
Let $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}$ be the basis of the integral lattice $\Gamma$ in Definition 1.3, let $I_{j}$ $(j=1,\ldots,n)$ be intervals of length $<1$, and let $E$ be the image of the set $\{\sum_{j=1}^{n}t_{j}\mathbf{e}_{j}:\,t_{j}\in I_{j}\}$ under the composite
$\mathfrak{h}_{\mathbf{R}}\to\mathfrak{h}_{\mathbf{R}}/\sum_{j=1}^{n}\mathbf{Z}\mathbf{e}_{j}\underset{\cong}{\stackrel{{\scriptstyle\exp(2\pi i\hbox{-})}}{{\to}}}T.$
For any $\boldsymbol{\lambda}\in I^{*}$, we have
$\int_{E}e(\boldsymbol{\lambda})\mathrm{d}\mu_{T}\ll\frac{1}{\mathbf{N}(\boldsymbol{\lambda})},$
where the constant implied by $\ll$ depends only on $n$.
Proof.
It suffices to treat the case where $\boldsymbol{\lambda}\not=0$. Let $\{\mathbf{e}^{*}_{1},\ldots,\mathbf{e}^{*}_{n}\}$ be the dual basis for $\mathfrak{h}^{*}_{\mathbf{R}}$, and write
$\boldsymbol{\lambda}=\sum_{j=1}^{n}\check{\lambda}_{j}\mathbf{e}^{*}_{j}.$
We have
$\int_{E}e(\boldsymbol{\lambda})\mathrm{d}\mu_{T}$ $=\int_{I_{1}\times\cdots\times I_{n}}e^{2\pi i\sum_{j=1}^{n}\check{\lambda}_{j}t_{j}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}.$
Let $\nabla$ be the gradient in the coordinate $(t_{1},\ldots,t_{n})$, let $\Delta=\nabla\cdot\nabla$ be the Laplace operator, and let $\|\boldsymbol{\lambda}\|^{\prime}=\sqrt{\sum_{j=1}^{n}\check{\lambda}_{j}^{2}}.$ We have
$\Delta\Big{(}e^{2\pi i\sum_{j=1}^{n}\check{\lambda}_{j}t_{j}}\Big{)}=-4\pi^{2}\|\boldsymbol{\lambda}\|^{\prime 2}e^{2\pi i\sum_{j=1}^{n}\check{\lambda}_{j}t_{j}}.$
By the Gauss formula, we have
$\int_{I_{1}\times\cdots\times I_{n}}e^{2\pi i\sum_{j=1}^{n}\check{\lambda}_{j}t_{j}}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}$
$=$ $-\frac{1}{4\pi^{2}\|\boldsymbol{\lambda}\|^{\prime 2}}\int_{I_{1}\times\cdots\times I_{n}}\Delta\Big{(}e^{2\pi i\sum_{j=1}^{n}\check{\lambda}_{j}t_{j}}\Big{)}\mathrm{d}t_{1}\cdots\mathrm{d}t_{n}$
$=$ $-\frac{1}{4\pi^{2}\|\boldsymbol{\lambda}\|^{\prime 2}}\int_{\partial(I_{1}\times\cdots\times I_{n})}\nabla\Big{(}e^{2\pi i\sum_{j=1}^{n}\check{\lambda}_{j}t_{j}}\Big{)}\cdot\mathbf{n}\,\mathrm{d}S$
$=$ $-\frac{1}{4\pi^{2}\|\boldsymbol{\lambda}\|^{\prime 2}}\int_{\partial(I_{1}\times\cdots\times I_{n})}2\pi ie^{2\pi i\sum_{j=1}^{n}\check{\lambda}_{j}t_{j}}(\check{\lambda}_{1},\cdots,\check{\lambda}_{n})\cdot\mathbf{n}\,\mathrm{d}S$
$\ll$ $\frac{1}{2\pi\|\boldsymbol{\lambda}\|^{\prime}}\mathrm{Area}(\partial(I_{1}\times\cdots\times I_{n})).$
[…]
Lemma 2.6.
Let $D$ be a small box in $T/W$, and let $M$ be a positive integer. There exists two linear combinations
$S^{\pm}=\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}\\ 0\leq{\lambda}_{j}\leq M\end{subarray}}\tilde{S}^{\pm}({\boldsymbol{\lambda}})\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}$
with $\tilde{S}^{\pm}({\boldsymbol{\lambda}})$ being complex numbers such that the following conditions holds:
  1. (i)
    $S^{\pm}$ are real valued functions on $T/W$, and
    $S^{-}\leq\chi_{D}\leq S^{+}$
    everywhere on $T/W$, where $\chi_{D}$ is the characteristic function of $D$.
  2. (ii)
    $\tilde{S}^{\pm}(0)=\int_{D}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T/W}+O\Big{(}\frac{1}{M}\Big{)}.$
  3. (iii)
    For any nonzero ${\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C}$ with $0\leq\lambda_{j}\leq M$, we have
    $\tilde{S}^{\pm}({\boldsymbol{\lambda}})\ll\frac{1}{\mathbf{N}(\boldsymbol{\lambda})}.$
Moreover, the implied constants in (ii)-(iii) are independent of $D$.
Proof.
The lattice of integral forms $I^{*}$ can be identified with the group of characters of $T$. Let $R(G_{\mathbf{R}})$ be the Grothendieck ring of representations of $G_{\mathbf{R}}$. The homomorphism
$\mathrm{char}:R(G_{\mathbf{R}})\to\mathbf{Z}[I^{*}]$
sending each representation $\rho$ of $G_{\mathbf{R}}$ to the character of $\rho|_{T}$ induces an isomorphism
$R(G_{\mathbf{R}})\cong\mathbf{Z}[I^{*}]^{W}.$
For any $L^{2}$ function $f$ on $T$, we have the Fourier expansion
$f=\sum_{\boldsymbol{\lambda}\in I^{*}}\hat{f}(\boldsymbol{\lambda})e(\boldsymbol{\lambda}),$
where
$\hat{f}(\boldsymbol{\lambda})=\int_{T}fe(-\boldsymbol{\lambda})\mathrm{d}\mu_{T}.$
Choose intervals $I_{j}$ $(j=1,\ldots,n)$ of lengths $<1$ so that $D\subset T/W$ is the image of the set
$E=\Big{\{}\exp\Big{(}2\pi i\sum_{j=1}^{n}t_{j}\mathbf{e}_{j}\Big{)}:\,t_{j}\in I_{j}\Big{\}}$
under the projection $T\to T/W$, and $E\cap\sigma(E)=\emptyset$ for every nonidentity element $\sigma$ in $W$. Let $L=[M/c]$, where $c$ is the constant in Lemma 2.4. By [Co85, Theorem 1, estimate above (10)], there exist two trigonometric polynomials
(2.6.1) $T^{\pm}=\sum_{\boldsymbol{\lambda}\in I^{*},\,|\check{\lambda}_{j}|\leq L}\hat{T}^{\pm}(\boldsymbol{\lambda})e(\boldsymbol{\lambda})$
such that the following conditions hold:
  1. (1)
    $T^{\pm}$ are real valued, and
    $T^{-}\leq\chi_{E}\leq T^{+}$
    everywhere on $T$, where $\chi_{E}$ is the characteristic function of $E$.
  2. (2)
    For any $\boldsymbol{\lambda}\in I^{*}$, we have
    $\hat{T}^{\pm}(\boldsymbol{\lambda})=\hat{\chi}_{E}(\boldsymbol{\lambda})+O\Big{(}\frac{1}{L}\Big{)},$
    where $\hat{\chi}_{E}(\boldsymbol{\lambda})=\int_{E}e(-\boldsymbol{\lambda})\mathrm{d}\mu_{T}$ and the constant implied by $O$ depends only on $n$.
For any trigonometric polynomial $p=\sum_{k=1}^{r}a_{k}e(\boldsymbol{\lambda}_{k})$, we have
$\int_{T}T^{\pm}p\,\mathrm{d}\mu_{T}=\sum_{k=1}^{r}a_{k}\int_{T}T^{\pm}e(\boldsymbol{\lambda}_{k})\,\mathrm{d}\mu_{T}=\sum_{k=1}^{r}a_{k}\hat{T}^{\pm}(-\boldsymbol{\lambda}_{k})$
$=$ $\sum_{k=1}^{r}a_{k}\hat{\chi}_{E}(-\boldsymbol{\lambda}_{k})+O\Big{(}\frac{1}{M}\Big{)}=\sum_{k=1}^{r}a_{k}\int_{E}e(\boldsymbol{\lambda}_{k})\,\mathrm{d}\mu_{T}+O\Big{(}\frac{1}{M}\Big{)}$
$=$ $\int_{E}p\,\mathrm{d}\mu_{T}+O\Big{(}\frac{1}{M}\Big{)}.$
that is, we have
(2.6.2) $\int_{T}T^{\pm}p\,\mathrm{d}\mu_{T}=\int_{E}p\,\mathrm{d}\mu_{T}+O\Big{(}\frac{1}{M}\Big{)}.$
Set
$S^{\pm}=\sum_{\sigma\in W}\sigma(T^{\pm}).$
Then $S^{\pm}$ are functions of $T$ invariant under $W$. Hence they define continuous functions on $T/W$ which we denote also by $S^{\pm}$. The inverse image of $D$ under the projection $T\to T/W$ is $\bigsqcup_{\sigma\in W}\sigma(E)$. By the condition (1) satisfied by $T^{\pm}$, we have
$S^{-}\leq\chi_{D}\leq S^{+}$
as functions on $T/W$. We have $S^{\pm}\in\mathbf{R}[I^{*}]^{W}$. So we can expand $S^{\pm}$ as a finite linear combination of the orthonormal basis $\{\frac{A_{{\boldsymbol{\lambda}}+\boldsymbol{\rho}}}{A_{\boldsymbol{\rho}}}:\,{\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C}\}$ of $L^{2}\Big{(}T/W,\big{(}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\big{)}\mu_{T/W}\Big{)}$. We can write
[…]
where
$\tilde{S}^{\pm}({\boldsymbol{\lambda}})=\frac{1}{|W|}\int_{T}S^{\pm}\overline{\Big{(}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\Big{)}}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}.$
In the expansion (2.6.3), the summation goes over those $\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}$ with $0\leq\lambda_{j}\leq M$. This follows from Lemma 2.4, and the fact that in the expansion
$S^{\pm}=\sum_{\sigma\in W}\sigma(T^{\pm})=\sum_{\sigma\in W,\,\boldsymbol{\lambda}\in I^{*},\,|\check{\lambda}_{j}|\leq L}\hat{T}^{\pm}(\boldsymbol{\lambda})e(\sigma(\boldsymbol{\lambda})),$
the summation goes over those $\boldsymbol{\lambda}\in I^{*}$ with $|\check{\lambda}_{j}|\leq L$.
By [BtD85, 1.7 (iii)], we have
$A_{\boldsymbol{\rho}}=e({\boldsymbol{\rho}})\prod_{{\boldsymbol{\alpha}}\in R^{+}}(1-e(-{\boldsymbol{\alpha}})).$
So we have
$A_{\boldsymbol{\rho}}\overline{A_{\boldsymbol{\rho}}}=\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}})),\quad\overline{\Big{(}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\Big{)}}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))=\overline{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}A_{\boldsymbol{\rho}}.$
Using the fact that $\overline{\Big{(}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\Big{)}}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}$ is $W$-invariant on $T$, we calculate $\tilde{S}^{\pm}({\boldsymbol{\lambda}})$ as follows:
$\tilde{S}^{\pm}({\boldsymbol{\lambda}})$ $=\frac{1}{|W|}\int_{T}S^{\pm}\overline{\Big{(}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\Big{)}}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}$
$=\frac{1}{|W|}\sum_{\sigma\in W}\int_{T}\sigma(T^{\pm})\overline{\Big{(}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\Big{)}}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}$
$=\int_{T}T^{\pm}\overline{\Big{(}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\Big{)}}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}$
$=\int_{T}T^{\pm}\overline{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}A_{\boldsymbol{\rho}}\,\mathrm{d}\mu_{T}$
$=\int_{E}\overline{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}A_{\boldsymbol{\rho}}\,\mathrm{d}\mu_{T}+O\Big{(}\frac{1}{M}\Big{)},$
where the last equality follows from equation (2.6.2) applied to the trigonometric polynomial $p=\overline{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}A_{\boldsymbol{\rho}}$. For ${\boldsymbol{\lambda}}=0$, we get
$\tilde{S}^{\pm}(0)$ $=\int_{E}\overline{A_{{\boldsymbol{\rho}}}}A_{\boldsymbol{\rho}}\,\mathrm{d}\mu_{T}+O\Big{(}\frac{1}{M}\Big{)}$
$=\int_{E}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}+O\Big{(}\frac{1}{M}\Big{)}$
$=\frac{1}{|W|}\int_{D}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T/W}+O\Big{(}\frac{1}{M}\Big{)}.$
For any nonzero ${\boldsymbol{\lambda}}$, we have
$\tilde{S}^{\pm}({\boldsymbol{\lambda}})$ $=\int_{E}\overline{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}A_{\boldsymbol{\rho}}\,\mathrm{d}\mu_{T}+O\Big{(}\frac{1}{M}\Big{)}$
$=\int_{E}\sum_{\sigma,\tau\in W}\mathrm{sgn}(\sigma\tau)e(\tau({\boldsymbol{\rho}})-\sigma(\boldsymbol{\rho})-\sigma({\boldsymbol{\lambda}}))\,\mathrm{d}\mu_{T}+O\Big{(}\frac{1}{M}\Big{)}$
$\ll\sum_{\sigma,\tau\in W}\frac{1}{\mathbf{N}(\tau({\boldsymbol{\rho}})-\sigma(\boldsymbol{\rho})-\sigma({\boldsymbol{\lambda}}))}+\frac{1}{M}$
$\ll\frac{1}{\mathbf{N}(\boldsymbol{\lambda})},$
where the first inequality follows from Lemma 2.5, and the last inequality follows from Lemma 2.3 and the fact that $0\leq\lambda_{j}\leq M$. ∎
Remark 2.7.
According to the constructions of Cochrane [Co85, Page 699] for $T^{\pm}$, as well as more precise evaluations in [BMV00], one can show that the trigonometric polynomials $T^{\pm}$ in (2.6.1) are uniformly bounded with bound depending only on $n$. It follows that $S^{\pm}$ are uniformly bounded with bound independent of $M$ and $E$.
2.8. Proof of Theorem 1.5
By Lemma 2.6 (i), we have
|$\frac{|\{1\leq i\leq N:x_{i}\in D\}|}{N}-\mu_{G_{\mathbf{R}}^{\natural}}(D)=\frac{1}{N}\sum_{i=1}^{N}\chi_{D}(x_{i})-\mu_{G_{\mathbf{R}}^{\natural}}(D)$
$\leq$ $\frac{1}{N}\sum_{i=1}^{N}S^{+}(x_{i})-\mu_{G_{\mathbf{R}}^{\natural}}(D)$
$=$ $\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{0\}\\ 0\leq\lambda_{j}\leq M\end{subarray}}\tilde{S}^{+}({\boldsymbol{\lambda}})\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})+\tilde{S}^{+}(0)-\mu_{G_{\mathbf{R}}^{\natural}}(D).$
Similarly, we have
|$\frac{|\{1\leq i\leq N:x_{i}\in D\}|}{N}-\mu_{G_{\mathbf{R}}^{\natural}}(D)$
$\geq$ $\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{0\}\\ 0\leq\lambda_{j}\leq M\end{subarray}}\tilde{S}^{-}({\boldsymbol{\lambda}})\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})+\tilde{S}^{-}(0)-\mu_{G_{\mathbf{R}}^{\natural}}(D).$
By Lemma 2.6 (ii)-(iii), we have
|$\frac{|\{1\leq i\leq N:x_{i}\in D\}|}{N}-\mu_{G_{\mathbf{R}}^{\natural}}(D)\ll\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{0\}\\ 0\leq\lambda_{j}\leq M\end{subarray}}\frac{1}{\mathbf{N}(\boldsymbol{\lambda})}\Big{|}\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})\Big{|}+\frac{1}{M}.$
3. Proof of Theorem 1.2
Throughout this section, $X$ is a smooth geometrically connected algebraic curve over $\mathbf{F}_{q}$, $\overline{X}$ a smooth compactification of $X$, $g$ the genus of $\overline{X}$, and $\eta$ the generic point of $X$. Any lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf $\mathcal{G}$ on $X$ corresponds to a representation $\rho_{\mathcal{G}}:\pi_{1}(X,\bar{\eta})\to\mathrm{GL}(\mathcal{G}_{\bar{\eta}})$. For any $x\in\overline{X}-X$, let $I_{x}$ be the inertia subgroup of $\pi_{1}(X,\bar{\eta})$ at $x$, and let $\mathrm{sw}_{x}(\mathcal{G})$ be the Swan conductor of $\rho_{\mathcal{G}}|_{I_{x}}$. We say $\mathcal{G}$ has no geometric invariant (resp. no geometric coinvariant) if
$\mathcal{G}_{\bar{\eta}}^{\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})}=0\quad(\hbox{resp. }\mathcal{G}_{\bar{\eta},\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})}=0).$
Proposition 3.1.
Let $\mathcal{G}$ be a lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$ punctually $\iota$-pure of weight 00. Suppose $\mathcal{G}$ has no geometric invariant. Let $N=|(\overline{X}-X)\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q}|$ and let $B$ be the largest break of $\rho_{\mathcal{G}}|_{I_{x}}$ $(x\in\overline{X}-X)$. We have
$\Big{|}\small{\sum_{x\in X(\mathbf{F}_{q^{m}})}\iota\mathrm{Tr}(F_{x},\mathcal{G}_{\bar{x}})}\Big{|}$ $\leq q^{m/2}(2g-2+N+NB)\mathrm{dim}(\mathcal{G}_{\bar{\eta}}).$
Proof.
By [De80, 1.4.1], we have
$H^{0}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})$$\cong\begin{cases}\mathcal{G}_{\bar{\eta}}^{\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})},&\text{if }X\text{ is proper},\\ 0,&\text{otherwise},\end{cases}$
$H^{2}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})$ $\cong\mathcal{G}_{\bar{\eta},\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})}(-1).$
Since $\mathcal{G}$ is punctually $\iota$-pure of weight 00, it is geometrically semisimple by [De80, 3.4.1(iii)]. Since $\mathcal{G}$ has no geometric invariant, it has no geometric coinvariant. Hence
$H^{0}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})=H^{2}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})=0.$
By the Grothendieck trace formula [De77, Rapport 3.2], we have
$\sum_{x\in X(\mathbf{F}_{q^{m}})}\mathrm{Tr}(F_{x},\mathcal{G}_{\bar{x}})=-\mathrm{Tr}\Big{(}F^{m},H^{1}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})\Big{)},$
where $F$ is the geometric Frobenius correspondence. By the Grothendieck–Ogg–Shafarevich formula [Gr77, X 7.12], we have
$\mathrm{dim}\,H^{1}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})$ $=(2g-2+N)\mathrm{dim}(\mathcal{G}_{\bar{\eta}})+\sum_{x\in X-U}\mathrm{deg}(x)\mathrm{sw}_{x}(\mathcal{G})$
$\leq(2g-2+N+NB)\mathrm{dim}(\mathcal{G}_{\bar{\eta}}).$
By Deligne’s theorem [De80, 3.3.1], all eigenvalues of $F$ on $H^{1}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})$ have absolute values $\leq q^{1/2}$ via $\iota$. Proposition 3.1 follows immediately. ∎
3.2. Proof of Theorem 1.2
For any nonzero $\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{0\}$, let
$\Gamma_{\boldsymbol{\lambda}}:G\to\mathrm{GL}(d,\overline{\mathbf{Q}}_{\ell})$
be the irreducible representation of $G$ with the highest weight $\boldsymbol{\lambda}$, where $d=\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})$. Let $\mathcal{G}_{\boldsymbol{\lambda}}$ be the $\overline{\mathbf{Q}}_{\ell}$-sheaf corresponding to the representation
$\pi_{1}(X,\bar{\eta})\stackrel{{\scriptstyle\rho_{\mathcal{F}}}}{{\to}}G\stackrel{{\scriptstyle\Gamma_{\boldsymbol{\lambda}}}}{{\to}}\mathrm{GL}(d,\overline{\mathbf{Q}}_{\ell}).$
Since $\Gamma_{\boldsymbol{\lambda}}$ is nontrivial and irreducible, the sheaf $\mathcal{G}_{\boldsymbol{\lambda}}$ has no geometric invariant. Moreover, $\mathcal{G}_{\boldsymbol{\lambda}}$ is punctually $\iota$-pure of weight 00 since $\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}$ lies in the maximal compact subgroup $G_{\mathbf{R}}$ of $G$ for every closed point $x$ in $X$. By the Weyl character formula, we have
$\iota\mathrm{Tr}(F_{x},\mathcal{G}_{\boldsymbol{\lambda},\bar{x}})=\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}).$
Applying Proposition 3.1 to the sheaf $\mathcal{G}_{\boldsymbol{\lambda}}$, we get
(3.2.1) $\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}})\Big{|}$ $=\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\iota\mathrm{Tr}(F_{x},\mathcal{G}_{\boldsymbol{\lambda},\bar{x}})\Big{|}$
$\leq\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})(2g-2+N+NB)q^{m/2},$
where $B$ is the largest break of of $\rho_{\mathcal{F}}|_{I_{x}}$ $(x\in\overline{X}-X)$. By the Weyl dimension formula [BtD85, VI 1.7 (iv)], we have
$\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})=\prod_{{\boldsymbol{\alpha}}\in R^{+}}\frac{({\boldsymbol{\lambda}}+{\boldsymbol{\rho}})(H_{\boldsymbol{\alpha}})}{{\boldsymbol{\rho}}(H_{\boldsymbol{\alpha}})}.$
It follows that
(3.2.2) $\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})\ll\mathbf{N}(\boldsymbol{\lambda})^{|R^{+}|}.$
Combining (3.2.1) and (3.2.2) with Theorem 1.5, for any positive integer $M$, we have
|$\ \ \ \ \frac{|\{x\in X(\mathbf{F}_{q^{m}}):\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D\}|}{|X(\mathbf{F}_{q^{m}})|}-\mu_{G_{\mathbf{R}}^{\natural}}(D)$
$\ll\frac{1}{|X(\mathbf{F}_{q^{m}})|}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{0\}\\ 0\leq\lambda_{j}\leq M\end{subarray}}\frac{1}{\mathbf{N}(\boldsymbol{\lambda})}\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}})\Big{|}+\frac{1}{M}$
$\ll\frac{1}{q^{m}}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{0\}\\ 0\leq\lambda_{j}\leq M\end{subarray}}\mathbf{N}(\boldsymbol{\lambda})^{|R^{+}|-1}q^{m/2}+\frac{1}{M}$
$\ll q^{-m/2}M^{|R^{+}|+n-1}+\frac{1}{M}.$
Taking $M$ to be the integer closest to $q^{\frac{m}{2(|R^{+}|+n)}}$, we get
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D\}|}{|X(\mathbf{F}_{q^{m}})|}=\mu_{G_{\mathbf{R}}^{\natural}}(D)+O\Big{(}q^{-\frac{m}{2(|R^{+}|+n)}}\Big{)}.$
4. Joint distribution
Denote data associated to $G^{\prime}$ by symbols with a superscript . For example, for each $\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}$, let $\lambda^{\prime}_{j}$ be the coefficients in the expression of $\boldsymbol{\lambda}^{\prime}$ as a linear combination of the fundamental weights for $G^{\prime}$.
Lemma 4.1.
Let $D,D^{\prime}$ be small boxes in $T/W$ and $T^{\prime}/W^{\prime}$, and let $M$ be a positive integer. There exist two trigonometric polynomial functions on $T/W\times T^{\prime}/W^{\prime}$ of the form
$P^{\pm}=\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ 0\leq{\lambda}_{j},\lambda^{\prime}_{j}\leq M\end{subarray}}c^{\pm}({\boldsymbol{\lambda}},{\boldsymbol{\lambda}^{\prime}})\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\frac{A_{{\boldsymbol{\lambda}^{\prime}}+{\boldsymbol{\rho}^{\prime}}}}{A_{\boldsymbol{\rho}^{\prime}}}+\Delta^{\pm}$
with $c^{\pm}({\boldsymbol{\lambda}},\boldsymbol{\lambda}^{\prime})$ being complex numbers and $\Delta^{\pm}$ being trigonometric polynomials such that the following conditions hold:
  1. (i)
    $P^{\pm}$ are real valued functions on $T/W\times T^{\prime}/W^{\prime}$, and
    $P^{-}\leq\chi_{D\times D^{\prime}}\leq P^{+}$
    everywhere on $T/W\times T^{\prime}/W^{\prime}$, where $\chi_{D\times D^{\prime}}$ is the the characteristic function of the box $D\times D^{\prime}$ in $T/W\times T^{\prime}/W^{\prime}$.
  2. (ii)
    We have
    $\Delta^{\pm}$ $\ll\frac{1}{M},$
    $c^{\pm}(0,0)$ $=\int_{D}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T/W}\cdot{\int_{D^{\prime}}}\frac{1}{|W^{\prime}|}\prod_{{\boldsymbol{\alpha}}\in R^{\prime}}(1-e({\boldsymbol{\alpha}^{\prime}}))\,\mathrm{d}\mu_{T^{\prime}/W^{\prime}}.$
  3. (iii)
    For any $\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}$ and $\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}$ such that $(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(0,0)$, we have
    $c^{\pm}({\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime}})\ll\frac{1}{\mathbf{N}(\boldsymbol{\lambda})\mathbf{N}(\boldsymbol{\lambda}^{\prime})}.$
Moreover, the implied constants in (ii)-(iii) are independent of $D$.
Proof.
Let $S^{\pm}=\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}\\ 0\leq{\lambda}_{j}\leq M\end{subarray}}\tilde{S}^{\pm}({\boldsymbol{\lambda}})\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}$ be the trigonometric polynomials constructed in Lemma 2.6. Let
$a^{\pm}(\boldsymbol{\lambda})$ $=$ $\begin{cases}\int_{D}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T/W}&\text{if }\boldsymbol{\lambda}=0,\\ \tilde{S}^{\pm}({\boldsymbol{\lambda}})\,&\text{if }\boldsymbol{\lambda}\not=0.\end{cases},$
$\delta^{\pm}$ $=$ $\tilde{S}^{\pm}(0)-\int_{D}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T/W}.$
Then
$S^{\pm}$ $=$ $\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}\\ 0\leq{\lambda}_{j}\leq M\end{subarray}}a^{\pm}({\boldsymbol{\lambda}})\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}+\delta^{\pm},$
$a^{\pm}(0)$ $=$ $\int_{D}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T/W},$
$a^{\pm}(\boldsymbol{\lambda})$ $\ll$ $\frac{1}{\mathbf{N}(\boldsymbol{\lambda})},\quad\delta^{\pm}\ll\frac{1}{M}.$
Write the trigonometric polynomials $S^{\prime\pm}$ constructed in Lemma 2.6 for $G^{\prime}$ as
$S^{\prime\pm}$ $=$ $\sum_{\begin{subarray}{c}\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ 0\leq{\lambda^{\prime}}_{j}\leq M\end{subarray}}a^{\prime\pm}({\boldsymbol{\lambda}^{\prime}})\frac{A_{{\boldsymbol{\lambda}^{\prime}}+{\boldsymbol{\rho}^{\prime}}}}{A_{\boldsymbol{\rho}^{\prime}}}+\delta^{\prime\pm},$
$a^{\prime\pm}(0)$ $=$ $\int_{D^{\prime}}\frac{1}{|W^{\prime}|}\prod_{{\boldsymbol{\alpha}^{\prime}}\in R^{\prime}}(1-e({\boldsymbol{\alpha}^{\prime}}))\,\mathrm{d}\mu_{T^{\prime}/W^{\prime}},$
$a^{\prime\pm}(\boldsymbol{\lambda}^{\prime})$ $\ll$ $\frac{1}{\mathbf{N}(\boldsymbol{\lambda}^{\prime})},\quad\delta^{\prime\pm}\ll\frac{1}{M}.$
Take $P^{+}=S^{+}S^{\prime+}.$ We have $\chi_{D}\chi_{D^{\prime}}\leq P^{+}$ and
$P^{+}=\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ 0\leq{\lambda}_{j},\lambda^{\prime}_{j}\leq M\end{subarray}}a^{+}({\boldsymbol{\lambda}})a^{\prime+}({\boldsymbol{\lambda}^{\prime}})\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\frac{A_{{\boldsymbol{\lambda}^{\prime}}+{\boldsymbol{\rho}^{\prime}}}}{A_{\boldsymbol{\rho}^{\prime}}}+\Delta^{+}$
with $\Delta^{+}$ given by
$\Delta^{+}=S^{+}\delta^{\prime+}+\delta^{+}S^{\prime+}-\delta^{+}\delta^{\prime+}.$
We have $\Delta\ll\frac{1}{M}$ since $\delta^{+},\delta^{\prime+}\ll\frac{1}{M}$ and $S^{+},S^{\prime+}\ll 1$ (Remark 2.7).
We have $(\chi_{D}-S^{-})(\chi_{D^{\prime}}-S^{\prime-})\geqslant 0,$ which implies
$\chi_{D}\chi_{D^{\prime}}\geq\chi_{D}S^{\prime-}+S^{-}\chi_{D^{\prime}}-S^{-}S^{\prime-}.$
Note that
$\chi_{D}S^{\prime-}=\chi_{D}(S^{\prime-}-S^{\prime+})+\chi_{D}S^{\prime+}\geq S^{+}(S^{\prime-}-S^{\prime+})+S^{-}S^{\prime+},$
$S^{-}\chi_{D^{\prime}}=(S^{-}-S^{+})\chi_{D^{\prime}}+S^{+}\chi_{D^{\prime}}\geq(S^{-}-S^{+})S^{\prime+}+S^{+}S^{\prime-}.$
Hence
$\chi_{D}\chi_{D^{\prime}}\geq 2(S^{+}S^{\prime-}+S^{-}S^{\prime+}-S^{+}S^{\prime+})-S^{-}S^{\prime-}.$
Take $P^{-}=2(S^{+}S^{\prime-}+S^{-}S^{\prime+}-S^{+}S^{\prime+})-S^{-}S^{\prime-}.$ As shown above for $S^{+}S^{\prime+}$, all the products $S^{+}S^{\prime-}$, $S^{-}S^{\prime+}$, $S^{+}S^{\prime+}$ and $S^{-}S^{\prime-}$ can be written in the form
$\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ 0\leq{\lambda}_{j},\lambda^{\prime}_{j}\leq M\end{subarray}}a({\boldsymbol{\lambda}},{\boldsymbol{\lambda}^{\prime}})\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\frac{A_{{\boldsymbol{\lambda}^{\prime}}+{\boldsymbol{\rho}^{\prime}}}}{A_{\boldsymbol{\rho}^{\prime}}}+\Delta$
such that
$a(0,0)=\int_{D}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T/W}\cdot{\int_{D^{\prime}}}\frac{1}{|W^{\prime}|}\prod_{{\boldsymbol{\alpha}}\in R^{\prime}}(1-e({\boldsymbol{\alpha}^{\prime}}))\,\mathrm{d}\mu_{T^{\prime}/W^{\prime}},$
$a({\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime}})\ll\frac{1}{\mathbf{N}(\boldsymbol{\lambda})\mathbf{N}(\boldsymbol{\lambda}^{\prime})},\quad\Delta\ll\frac{1}{M}.$
It follows that $P^{-}$ has the required properties. ∎
Remark 4.2.
In the case where $G=G^{\prime}=\mathrm{SL}_{2}$, the result corresponding to Lemma 4.1 is proved in [FLLX24, Lemma 5.1]. We follow the same proof here.
Proposition 4.3.
For any sequences $x_{1},\ldots,x_{N}\in G_{\mathbf{R}}^{\natural}$ and $x^{\prime}_{1},\ldots,x^{\prime}_{N}\in G^{\prime\natural}_{\mathbf{R}}$, any small boxes $D$ in $G_{\mathbf{R}}^{\natural}$ and $D^{\prime}$ in $G^{\prime\natural}_{\mathbf{R}}$, and any positive integer $M$, we have
|$\frac{|\{1\leq i\leq N:x_{i}\in D,\,x^{\prime}_{i}\in D^{\prime}\}|}{N}-\mu_{G_{\mathbf{R}}^{\natural}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})$
$\ll\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(0,0)\\ 0\leq\lambda_{j},\lambda^{\prime}_{j}\leq M\end{subarray}}\frac{1}{\mathbf{N}(\boldsymbol{\lambda})\mathbf{N}(\boldsymbol{\lambda}^{\prime})}\Big{|}\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})\frac{A_{{\boldsymbol{\lambda}^{\prime}}+{\boldsymbol{\rho}^{\prime}}}}{A_{\boldsymbol{\rho}^{\prime}}}(x^{\prime}_{i})\Big{|}+\frac{1}{M},$
where the constant implied by $\ll$ depends only on $G$ and $G^{\prime}$.
Proof.
By Lemma 4.1 (i), we have
|$\frac{|\{1\leq i\leq N:x_{i}\in D,x^{\prime}_{i}\in D^{\prime}\}|}{N}-\mu_{G_{\mathbf{R}}^{\natural}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})$
$=$ $\frac{1}{N}\sum_{i=1}^{N}\chi_{D\times D^{\prime}}(x_{i},x^{\prime}_{i})-\mu_{G_{\mathbf{R}}^{\natural}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})$
$\leq$ $\frac{1}{N}\sum_{i=1}^{N}P^{+}(x_{i},x^{\prime}_{i})-\mu_{G_{\mathbf{R}}^{\natural}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})$
$=$ $\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(0,0)\\ 0\leq\lambda_{j},\lambda^{\prime}_{j}\leq M\end{subarray}}c^{+}(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})\frac{A_{{\boldsymbol{\lambda}^{\prime}}+{\boldsymbol{\rho}^{\prime}}}}{A_{\boldsymbol{\rho}^{\prime}}}(x^{\prime}_{i})+\Delta^{+}(x_{i},x^{\prime}_{i}).$
Similarly, we have
|$\frac{|\{1\leq i\leq N:x_{i}\in D,x^{\prime}_{i}\in D^{\prime}\}|}{N}-\mu_{G_{\mathbf{R}}^{\natural}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})$
$\geq$ $\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(0,0)\\ 0\leq\lambda_{j},\lambda^{\prime}_{j}\leq M\end{subarray}}c^{-}(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})\frac{A_{{\boldsymbol{\lambda}^{\prime}}+{\boldsymbol{\rho}^{\prime}}}}{A_{\boldsymbol{\rho}^{\prime}}}(x^{\prime}_{i})+\Delta^{-}(x_{i},x^{\prime}_{i}).$
By Lemma 4.1 (ii)-(iii), we have
|$\frac{|\{1\leq i\leq N:x_{i}\in D,x^{\prime}_{i}\in D^{\prime}\}|}{N}-\mu_{G_{\mathbf{R}}^{\natural}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})$
$\ll$ $\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(0,0)\\ 0\leq\lambda_{j},\lambda^{\prime}_{j}\leq M\end{subarray}}\frac{1}{\mathbf{N}(\boldsymbol{\lambda})\mathbf{N}(\boldsymbol{\lambda}^{\prime})}\Big{|}\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})\frac{A_{{\boldsymbol{\lambda}^{\prime}}+{\boldsymbol{\rho}^{\prime}}}}{A_{\boldsymbol{\rho}^{\prime}}}(x^{\prime}_{i})\Big{|}+\frac{1}{M}.$
4.4. Proof of Theorem 1.7
For any $\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}$, let $\Gamma_{\boldsymbol{\lambda}^{\prime}}$ be the irreducible representation of $G^{\prime}$ with the highest weight $\boldsymbol{\lambda}^{\prime}$, and let $\mathcal{G}_{\boldsymbol{\lambda}^{\prime}}$ be the $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$ corresponding to the representation $\Gamma_{\boldsymbol{\lambda}^{\prime}}\circ\rho_{\mathcal{F}^{\prime}}$. By our assumption, if $(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(0,0)$, the sheaf $\mathcal{G}_{\boldsymbol{\lambda}}\otimes\mathcal{G}_{\boldsymbol{\lambda}^{\prime}}$ has no geometric invariant and is punctually $\iota$-pure of weight 00. By the Weyl character formula, we have
$\iota\mathrm{Tr}\Big{(}F_{x},(\mathcal{G}_{\boldsymbol{\lambda}}\otimes\mathcal{G}_{\boldsymbol{\lambda}^{\prime}})_{\bar{x}}\Big{)}=\frac{A_{\boldsymbol{\lambda}+\boldsymbol{\rho}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}})\frac{A_{\boldsymbol{\lambda}^{\prime}+\boldsymbol{\rho}^{\prime}}}{A_{\boldsymbol{\rho}^{\prime}}}(\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{ss}}).$
Applying Proposition 3.1 to the sheaf $\mathcal{G}_{\boldsymbol{\lambda}}\otimes\mathcal{G}_{\boldsymbol{\lambda}^{\prime}}$, we get
$\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\frac{A_{\boldsymbol{\lambda}+\boldsymbol{\rho}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}})\frac{A_{\boldsymbol{\lambda}^{\prime}+\boldsymbol{\rho}^{\prime}}}{A_{\boldsymbol{\rho}^{\prime}}}(\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{ss}})\Big{|}\ll\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}^{\prime}})q^{\frac{m}{2}}.$
Note that
$\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}^{\prime}})\ll\mathbf{N}(\boldsymbol{\lambda})^{|R^{+}|}\cdot\mathbf{N}(\boldsymbol{\lambda}^{\prime})^{|R^{\prime+}|}.$
By Proposition 4.3, for any positive integer $M$, we have
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\,\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D,\,\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{s}s}\in D^{\prime}\}|}{|X(\mathbf{F}_{q^{m}})|}-\mu_{G_{\mathbf{R}}^{\natural}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})$
$\ll$ $\frac{1}{|X(\mathbf{F}_{q^{m}})|}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(0,0)\\ 0\leq\lambda_{j},\lambda^{\prime}_{j}\leq M\end{subarray}}\frac{1}{\mathbf{N}(\boldsymbol{\lambda})\mathbf{N}(\boldsymbol{\lambda}^{\prime})}$
×$\quad\times\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\frac{A_{\boldsymbol{\lambda}+\boldsymbol{\rho}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}})\frac{A_{\boldsymbol{\lambda}^{\prime}+\boldsymbol{\rho}^{\prime}}}{A_{\boldsymbol{\rho}^{\prime}}}(\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{ss}})\Big{|}+\frac{1}{M}$
$\ll$ $\frac{1}{q^{m}}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(0,0)\\ 0\leq\lambda_{j},\lambda^{\prime}_{j}\leq M\end{subarray}}\mathbf{N}(\boldsymbol{\lambda})^{|R^{+}|-1}\mathbf{N}(\boldsymbol{\lambda}^{\prime})^{|R^{\prime+}|-1}q^{m/2}+\frac{1}{M}$
$\ll$ $q^{-m/2}M^{|R^{+}|+|R^{\prime+}|+n+n^{\prime}-2}+\frac{1}{M}.$
Taking $M$ to be the integer closest to $q^{\frac{m}{2(|R^{+}|+|R^{\prime+}|+n+n^{\prime}-1)}}$, we get
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\,\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D,\,\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{s}s}\in D^{\prime}\}|}{|X(\mathbf{F}_{q^{m}})|}$
$=$ $\mu_{G^{\natural}_{\mathbf{R}}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})+O\Big{(}q^{-\frac{m}{2(|R^{+}|+|R^{\prime+}|+n+n^{\prime}-1)}}\Big{)}.$
This completes the proof of Theorem 1.7.
4.5. Joint distribution of the Kloosterman sums and the Airy sums.
For any $x\in\mathbb{G}_{m}(\mathbf{F}_{q^{m}})\cong\mathbf{F}^{*}_{q^{m}}$, the Kloosterman sum is defined by
$\mathrm{Kl}(\mathbf{F}_{q^{m}},x)=\sum_{z\in\mathbf{F}^{*}_{q^{m}}}e^{\frac{2\pi i}{p}\mathrm{Tr}_{\mathbf{F}_{q^{m}}/\mathbf{F}_{p}}(z+\frac{x}{z})}.$
Deligne [De77, Sommes trig. 7.8] constructs the Kloosterman sheaf $\mathrm{Kl}$ which is a lisse $\overline{\mathbf{Q}}_{\ell}$-adic sheaf on $\mathbb{G}_{m}$ punctually $\iota$-pure of weight $1$ such that
$\iota\mathrm{Tr}(F_{x},\mathrm{Kl}_{\bar{x}})=-\mathrm{Kl}(\mathbf{F}_{q^{m}},x).$
for all $x\in\mathbb{G}_{m}(\mathbf{F}_{q^{m}})$. Consider the twist $\mathrm{Kl}(1/2)$ of $\mathrm{Kl}$ by $\mathbf{Q}_{\ell}(1/2)$ so that $\mathrm{Kl}(1/2)$ is punctually $\iota$-pure of weight 00. Katz ([Ka88, 11.1, 11.3]) shows that both the arithmetic and the geometric monodromy groups of $\mathrm{Kl}(1/2)$ are $\mathrm{SL}_{2}$. Define an angle $\theta(x)$ by
$\mathrm{Kl}(\mathbf{F}_{q^{m}},x)$ $=2q^{m/2}\cos\theta(x).$
For any $x\in\mathbb{A}^{1}(\mathbf{F}_{q^{m}})\cong\mathbf{F}_{q^{m}}$, the Airy sum is defined by
$\mathrm{Ai}(\mathbf{F}_{q^{m}},x)=\sum_{z\in\mathbf{F}_{q^{m}}}e^{\frac{2\pi i}{p}\mathrm{Tr}_{\mathbf{F}_{q^{m}}/\mathbf{F}_{p}}(z^{3}+xz)}.$
Katz ([Ka87, Theorem 17]) constructs the Airy sheaf $\mathrm{Ai}$ which is a lisse $\overline{\mathbf{Q}}_{\ell}$-adic sheaf on $\mathbb{A}^{1}$ punctually pure of weight $1$ such that
$\iota\mathrm{Tr}(F_{x},\mathrm{Ai}_{\bar{x}})=-\mathrm{Ai}(\mathbf{F}_{q^{m}},x).$
for all $x\in\mathbb{A}^{1}(\mathbf{F}_{q^{m}})$. Katz ([Ka87, Theorem 17 (iii), Corollary 20]) proves that $\mathrm{det}(\mathrm{Ai})$ is geometrically constant, and is given by the character $\mathrm{Gal}(\overline{\mathbf{F}}_{q}/\mathbf{F}_{q})\to\overline{\mathbf{Q}}_{\ell}^{*}$ mapping the geometric Frobenius element in $\mathrm{Gal}(\overline{\mathbf{F}}_{q}/\mathbf{F}_{q})$ to $\prod_{\chi^{3}=1,\,\chi\not=1}(-g(\chi,\psi))$, where the product is taken over all nontrivial multiplicative characters $\chi:\mathbf{F}^{*}_{q}\to\overline{\mathbb{Q}}_{\ell}^{*}$ of order dividing $3$ and
$g(\chi,\psi)=\sum_{x\in\mathbf{F}^{*}_{q}}\chi(x)\psi(x)$
is the Gauss sum. Let $\alpha$ be a square root of $\prod_{\chi^{3}=1,\,\chi\not=1}(-g(\chi,\psi))^{-1}$. Let $\mathrm{Ai}(\alpha)$ be the twist of the Airy sheaf by the character $\mathrm{Gal}(\overline{\mathbf{F}}_{q}/\mathbf{F}_{q})\to\overline{\mathbf{Q}}_{\ell}^{*}$ sending the geometric Frobenius element in $\mathrm{Gal}(\overline{\mathbf{F}}_{q}/\mathbf{F}_{q})$ to $\alpha$. Then $\mathrm{Ai}(\alpha)$ is punctually $\iota$-pure of weight 00 and $\mathrm{det}(\mathrm{Ai}(\alpha))=1$. Katz ([Ka87, Theorem 19]) shows that both the arithmetic and the geometric monodromy group of $\mathrm{Ai}(\alpha)$ is $\mathrm{SL}_{2}$ if $p>7$. Define an angle $\theta^{\prime}(x)$ by
$\mathrm{Ai}(\mathbf{F}_{q^{m}},x)$ $=2\alpha^{-m}\cos\theta^{\prime}(x).$
Corollary 4.6.
Suppose $p>7$. For any subintervals $I,I^{\prime}\subset[0,\pi],$ we have
$\frac{1}{q^{m}-1}\Big{|}\{x\in\mathbb{G}_{m}(\mathbf{F}_{q^{m}}):\theta(x)\in I,\;\theta^{\prime}(x)\in I^{\prime}\}\Big{|}=\mu_{\mathrm{ST}}(I)\mu_{\mathrm{ST}}(I^{\prime})+O(q^{-\frac{m}{6}}),$
where $\mu_{ST}=\frac{2}{\pi}\sin^{2}\theta\,\mathrm{d}\theta$ is the Sato-Tate measure on $[0,\pi]$.
Proof.
We apply Theorem 1.7 to the case where $\mathcal{F}=\mathrm{Kl}(1/2)$ and $\mathcal{F}^{\prime}=\mathrm{Ai}(\alpha)$. For $\mathfrak{sl}_{2}$, the dimension of a Cartan subalgebra is
$n=n^{\prime}=1.$
The number of positive roots is also 1. So we have
$62(|R^{+}|+|R^{\prime+}|+n+n^{\prime}-1)=$ 6 .
A maximal compact subgroup of $\mathrm{SL}_{2}(\mathbf{C})$ is $\mathrm{SU}(2)$, and any element in $\mathrm{SU}(2)$ is conjugate to $(\begin{smallmatrix}e^{i\theta}&\\ &e^{-i\theta}\end{smallmatrix})$ for a unique $\theta\in[0,\pi]$. So $\mathrm{SU}(2)^{\natural}$ is identified with $[0,\pi]$. For any $x\in\mathbb{G}_{m}(\mathbf{F}_{q^{m}})$, the conjugacy class $\rho_{\mathrm{Kl}(1/2)}(F_{x})^{\mathrm{ss}}\in\mathrm{SU}(2)^{\natural}$ corresponds to the angle $\theta(x)\in[0,\pi]$, and $\rho_{\mathrm{Ai}(\alpha)}(F_{x})^{\mathrm{ss}}\in\mathrm{SU}(2)^{\natural}$ corresponds to the angle $\theta^{\prime}(x)\in[0,\pi]$. Using Weyl’s integration formula, one can verify $\mu_{\mathrm{SU}(2)^{\natural}}$ is identified with the Sato-Tate measure on $[0,\pi]$. Confer Proposition 2.1 below. Irreducible representations of $\mathrm{SL}_{2}$ are symmetric products of the standard representation. We claim that for any positive integers $k_{1},k_{2}$, $\mathrm{Sym}^{k_{1}}(\mathrm{Kl})\otimes\mathrm{Sym}^{k_{2}}(\mathrm{Ai})$ has no nonzero $\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})$-invariant. Our assertion then follows immediately from Theorem 1.7.
By [Ka88, 4.1.3 and 4.1.4], $\mathrm{Kl}$ is geometrically self dual. So we have
$\ \ \ \ \Big{(}\mathrm{Sym}^{k_{1}}(\mathrm{Kl})_{\bar{\eta}}\otimes\mathrm{Sym}^{k_{2}}(\mathrm{Ai})_{\bar{\eta}}\Big{)}^{\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})}$
$\cong\Big{(}\mathrm{Sym}^{k_{1}}(\mathrm{Kl})_{\bar{\eta}}^{\vee}\otimes\mathrm{Sym}^{k_{2}}(\mathrm{Ai})_{\bar{\eta}}\Big{)}^{\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})}$
$\cong\mathrm{Hom}_{\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})}\Big{(}\mathrm{Sym}^{k_{1}}(\mathrm{Kl})_{\bar{\eta}},\mathrm{Sym}^{k_{2}}(\mathrm{Ai})_{\bar{\eta}}\Big{)}.$
By [De77, Sommes trig. 7.8 (iii)], the sheaf $\mathrm{Sym}^{k_{1}}(\mathrm{Kl})$ is ramified at 00 for all $k_{1}\geq 1$. But $\mathrm{Sym}^{k_{2}}(\mathrm{Ai})$ is unramified at 00. So $\mathrm{Sym}^{k_{1}}(\mathrm{Kl})_{\bar{\eta}}$ and $\mathrm{Sym}^{k_{2}}(\mathrm{Ai})_{\bar{\eta}}$ are non-isomorphic irreducible representations of $\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})$. By Schur’s lemma, we have
$\mathrm{Hom}_{\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})}\Big{(}\mathrm{Sym}^{k_{1}}(\mathrm{Kl})_{\bar{\eta}},\mathrm{Sym}^{k_{2}}(\mathrm{Ai})_{\bar{\eta}}\Big{)}=0.$
So $\mathrm{Sym}^{k_{1}}(\mathrm{Kl})\otimes\mathrm{Sym}^{k_{2}}(\mathrm{Kl})$ has no nonzero $\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})$-invariant. ∎