### v2

# An Effective Deligne’s equidistribution theorem

Abstract.
We prove an Erdős–Turán type inequality for compact Lie groups, from which we deduce an effective version of Deligne’s equidistribution theorem.
Key words and phrases:
Deligne’s equidistribution theorem, Erdős–Turán inequality, Weyl integration formula, Weyl character formula.
2020 Mathematics Subject Classification:
11K38, 22E46, 14F20
1. Introduction
Let $\mathbf{F}_{q}$ be a finite field of characteristic $p$ with $q$ elements, $X$ a smooth geometrically connected scheme over $\mathbf{F}_{q}$, $\eta$ the generic point of $X$, $\ell$ a prime number distinct from $p$, and $\mathcal{F}$ a lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$. We have an exact sequence of fundamental groups
$1\to\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})\to\pi_{1}(X,\bar{\eta})\to\mathrm{Gal}(\overline{\mathbf{F}}_{q}/\mathbf{F}_{q})\to 1$
The lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf $\mathcal{F}$ on $X$ defines a representation
$\rho_{\mathcal{F}}:\pi_{1}(X,\bar{\eta})\to\mathrm{GL}(\mathcal{F}_{\bar{\eta}}).$
The geometric monodromy group (resp. arithmetic monodromy group) of $\mathcal{F}$ is defined to be the Zariski closure of the image of $\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})$ (resp. $\pi_{1}(X,\bar{\eta}))$ in $\mathrm{GL}(\mathcal{F}_{\bar{\eta}})$.
Fix an isomorphism $\iota:\overline{\mathbf{Q}}_{\ell}\stackrel{{\scriptstyle\cong}}{{\to}}\mathbf{C}$. Throughout this paper, we assume $\mathcal{F}$ is punctually $\iota$-pure of weight 00, that is, for any closed point $x$ in $X$ and any eigenvalue $\lambda$ of the action of the geometric Frobenius element $F_{x}$ on $\mathcal{F}_{\bar{x}}$, we have $|\iota(\lambda)|=1$. Then by theorems of Deligne and Grothendieck [De80, 3.4.1(iii), 1.3.9], the identity component $G^{0}$ of the geometric monodromy group $G$ is semisimple. Via the isomorphism $\iota$, the algebraic group $G$ defines a complex semisimple Lie group $G_{\mathbf{C}}$. The category of finite dimensional $\overline{\mathbf{Q}}_{\ell}$-representations of $G$ is equivalent to that of finite dimensional complex representations of $G_{\mathbf{C}}$.
We assume $G$ is connected throughout this paper. Let $\mathfrak{g}$ be the Lie algebra of $G_{\mathbf{C}}$. Fix a Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{g}$, and let $R$ be the root system. Choose an ordering $R=R^{+}\cup R^{-}$ on $R$. For each ${\boldsymbol{\alpha}}\in R$, let $H_{\boldsymbol{\alpha}}\in\mathfrak{h}$ be the element in $[\mathfrak{g}_{\boldsymbol{\alpha}},\mathfrak{g}_{-{\boldsymbol{\alpha}}}]$ satisfying ${\boldsymbol{\alpha}}(H_{\boldsymbol{\alpha}})=2$, and let $\mathfrak{h}_{\mathbf{R}}$ be the real subspace of $\mathfrak{h}$ spanned by $H_{\boldsymbol{\alpha}}$ $(\boldsymbol{\alpha}\in R)$. There exists a real subalgebra $\mathfrak{g}_{\mathbf{R}}$ of $\mathfrak{g}$ such that
$\mathfrak{g}_{\mathbf{R}}=i\mathfrak{h}_{\mathbf{R}}\bigoplus\bigoplus_{{\boldsymbol{\alpha}}\in R^{+}}\Big{(}\mathfrak{g}_{\mathbf{R}}\cap(\mathfrak{g}_{\boldsymbol{\alpha}}\oplus\mathfrak{g}_{-{\boldsymbol{\alpha}}})\Big{)},$
that $\mathfrak{g}_{\mathbf{R}}$ is the Lie algebra of a maximal compact subgroup ${G_{\mathbf{R}}}$ of $G$, and that $i\mathfrak{h}_{\mathbf{R}}$ is the Lie algebra of a maximal torus $T$ of ${G_{\mathbf{R}}}$. The category of finite dimensional complex representations of $G_{\mathbf{C}}$ is equivalent to that of finite dimensional complex representations of $G_{\mathbf{R}}$.
For simplicity, we assume $\rho_{\mathcal{F}}(\pi_{1}(X,\bar{\eta}))\subset G$ as in [Ka88, 3.3], that is, the arithmetic monodromy group coincides with the geometric monodromy group. Then for every closed point $x$ in $X$, the geometric Frobenius conjugacy class $F_{x}$ in $\pi_{1}(U,\bar{\eta})^{\natural}$ defines an element in $G^{\natural}$ and an element in $G^{\natural}_{\mathbf{C}}$, where the set of conjugacy classes of a group $H$ is denoted by $H^{\natural}$. Since we assume $\mathcal{F}$ is punctually $\iota$-pure of weight 00, the semisimple part of $\rho_{\mathcal{F}}(F_{x})$ defines an element in $G_{\mathbf{R}}^{\natural}$ which we denote by $\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}$. Provide $G_{\mathbf{R}}^{\natural}$ with the quotient topology induced from $G_{\mathbf{R}}$, and let $\mu_{G_{\mathbf{R}}^{\natural}}$ be the measure on $G_{\mathbf{R}}^{\natural}$ induced by the Haar measure $\mu_{G_{\mathbf{R}}}$ on $G_{\mathbf{R}}$. For any integrable function $f$ on $G_{\mathbf{R}}^{\natural}$, we have
$\int_{G_{\mathbf{R}}^{\natural}}f\,\mathrm{d}\mu_{G_{\mathbf{R}}^{\natural}}=\int_{G_{\mathbf{R}}}f\,\mathrm{d}\mu_{G_{\mathbf{R}}}.$
Deligne proves the following theorem ([De80, 3.5.3], [Ka88, 3.6]).
Theorem 1.1 (Deligne).
Let $\mathcal{F}$ be a lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$ punctually $\iota$-pure of weight 00, and let $G$ be its geometric monodromy group. Suppose the arithmetic monodromy group of $\mathcal{F}$ coincides with $G$. For any positive integer $m$, let $X(\mathbf{F}_{q^{m}})$ be the set of $\mathbf{F}_{q^{m}}$-points in $X$. For any domain $D$ in $G_{\mathbf{R}}^{\natural}$ so that the boundary of $D$ has measure 00, we have
$\lim_{m\to\infty}\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D\}|}{|X(\mathbf{F}_{q^{m}})|}=\mu_{G_{\mathbf{R}}^{\natural}}(D).$
The main result of this paper is an effective version of the above theorem.
Theorem 1.2.
Let $\mathcal{F}$ be a lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$ punctually $\iota$-pure of weight 00, and let $G$ be its geometric monodromy group. Suppose $G$ is connected and the arithmetic monodromy group of $\mathcal{F}$ coincides with $G$. Then for any small box $D$ in $G^{\natural}_{\mathbf{R}}$, we have
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D\}|}{|X(\mathbf{F}_{q^{m}})|}=\mu_{G_{\mathbf{R}}^{\natural}}(D)+O\Big{(}q^{-\frac{m}{2(|R^{+}|+1)}}(\log q^{m})^{\frac{n-1}{|R^{+}|+1}}\Big{)},$
where $n:=\mathrm{dim}_{\mathbf{C}}\mathfrak{h}$ is the rank of the Lie group, and the constant implied by $O$ depends only on $X$ and $\mathcal{F}$.
We will describe small boxes in Definition 1.3 below. In the case where $X$ is an algebraic curve, let $\overline{X}$ be a smooth compactification of $X$. Then the constant implied by $O$ in the theorem depends only on the group $G$, the genus of $\overline{X}$, the number of points in $(\overline{X}-X)(\overline{\mathbf{F}}_{q})$, the rank of $\mathcal{F}$ and the Swan conductors of $\mathcal{F}|_{I_{x}}$ ($x\in\overline{X}-X$), where $I_{x}$ is the inertia subgroup of $\pi_{1}(X,\bar{\eta})$ at $x$.
We deduce the above theorem from an Erdős-Turán type inequality for compact Lie groups, which is of independent interest and we now describe. Let
$N(T):=\{g\in{G_{\mathbf{R}}}:gTg^{-1}=T\}$
be the normalizer of the maximal torus $T$ in ${G_{\mathbf{R}}}$, and let $W:=N(T)/T$ be the Weyl group. It acts on $T$ by
$W\times T\to T,\quad(gT,x)\mapsto gxg^{-1}.$
Hence $W$ acts on the Lie algebra $\mathfrak{h}_{\mathbf{R}}$ and its dual $\mathfrak{h}_{\mathbf{R}}^{*}$. Provide $\mathfrak{h}_{\mathbf{R}}$ and $\mathfrak{h}_{\mathbf{R}}^{*}$ with the inner product induced by the Killing form. The Weyl group can be identified with the subgroup of $\mathrm{GL}(\mathfrak{h}_{\mathbf{R}}^{*})$ generated by reflections with respect to ${\boldsymbol{\alpha}}^{\perp}$ $({\boldsymbol{\alpha}}\in R)$. Let $T/W$ be the quotient space. The map $T\to G_{\mathbf{R}}^{\natural}$ sending each element in $T$ to its conjugacy class in $G_{\mathbf{R}}$ defines a homeomorphism
$\kappa:T/W\stackrel{{\scriptstyle\cong}}{{\to}}G_{\mathbf{R}}^{\natural}$
by [BtD85, IV Proposition 2.6] if we provide $T/W$ and $G_{\mathbf{R}}^{\natural}$ with the quotient topology induced from $T$ and $G_{\mathbf{R}}$, respectively.
The map $\exp:2\pi i\mathfrak{h}_{\mathbf{R}}\to T$ is an epimormphism. Let $2\pi i\Gamma$ be its kernel. Then we have an isomorphism
$\exp(2\pi i\hbox{-}):\mathfrak{h}_{\mathbf{R}}/\Gamma\stackrel{{\scriptstyle\cong}}{{\to}}T.$
We call $\Gamma$ the integral lattice. Define the lattice of integral forms $I^{*}$ to be
$I^{*}:=\{{\boldsymbol{\lambda}}\in\mathfrak{h}_{\mathbf{R}}^{*}:\,{\boldsymbol{\lambda}}(\Gamma)\subset\mathbf{Z}\}.$
For any ${\boldsymbol{\lambda}}\in I^{*}$, define a character
$e({\boldsymbol{\lambda}}):T\to S^{1}:=\{z\in\mathbf{C}:\,|z|=1\}$
which maps $\exp(2\pi iH)\in T$ to $e^{2\pi i{\boldsymbol{\lambda}}(H)}\in S^{1}$ for every $H\in\mathfrak{h}_{\mathbf{R}}$. We thus get an isomorphism
$I^{*}\stackrel{{\scriptstyle\cong}}{{\to}}\mathrm{Hom}(T,S^{1}),\quad\boldsymbol{\lambda}\mapsto e(\boldsymbol{\lambda}).$
For any weight ${\boldsymbol{\lambda}}\in\mathfrak{h}^{*}$ of a representation of $\mathfrak{g}$ coming from a representation of $G$, we have $\boldsymbol{\lambda}\in I^{*}$. In particular, for any root ${\boldsymbol{\alpha}}\in R$, we have $\alpha\in I^{*}$ and we have a character $e({\boldsymbol{\alpha}})$ of $T$. Define the weight lattice to be
$\Lambda:=\{{\boldsymbol{\lambda}}\in\mathfrak{h}_{\mathbf{R}}^{*}:\,{\boldsymbol{\lambda}}(H_{\boldsymbol{\alpha}})\in\mathbf{Z}\hbox{ for all }{\boldsymbol{\alpha}}\in R\}.$
We have
$\sum_{\boldsymbol{\alpha}\in R}\mathbf{Z}H_{\boldsymbol{\alpha}}~{}\subseteq$ $~{}\Gamma~{}\subseteq~{}\{H\in\mathfrak{h}_{\mathbf{R}}:\,\boldsymbol{\alpha}(H)\in\mathbf{Z}\hbox{ for all }\boldsymbol{\alpha}\in R\},$
$\Lambda~{}\supseteq$ $~{}I^{*}~{}\supseteq~{}\sum_{\boldsymbol{\alpha}\in R}\mathbf{Z}{\boldsymbol{\alpha}}.$
Definition 1.3.
Fix a basis $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}$ for the lattice $\Gamma$. A subset of $T$ of the form
$\Omega=\Big{\{}\exp\Big{(}2\pi i\sum_{j=1}^{n}t_{j}\mathbf{e}_{j}\Big{)}:\,t_{j}\in I_{j}\Big{\}}$
for some intervals $I_{j}$ of lengths $<1$ is called a box in $T$. We say the box $\Omega$ is small if for any nonidentity element $\sigma$ in the Weyl group $W$, $\Omega\cap\sigma(\Omega)$ is empty. A subset $D$ of $T/W$ (resp. $G_{\mathbf{R}}^{\natural}$) is called a small box if it is the image of a small box $\Omega$ in $T$ under the canonical map
$T\to T/W\quad(\mathrm{resp.}~{}T\to T/W\underset{\cong}{\stackrel{{\scriptstyle\kappa}}{{\to}}}G_{\mathbf{R}}^{\natural}).$
Remark 1.4.
With the above notations, since the lengths of the intervals $I_{j}$ are $<1$, the map $\exp(2\pi i\hbox{-}):\mathfrak{h}_{R}\to T$ identifies the set $I_{1}\times\cdots\times I_{n}$ with its image $\Omega$ in $T$. Suppose furthermore that $\Omega\cap\sigma(\Omega)$ is empty for every nonidentity element $\sigma\in W$. Then the projection $T\to T/W$ identifies $\Omega$ with its image $D$ in $T/W$, and the inverse image of $D$ in $T$ is $\bigsqcup_{\sigma\in W}\sigma(\Omega)$.
Let
:=$\mathfrak{C}:=\{{\boldsymbol{\lambda}}\in\mathfrak{h}_{\mathbf{R}}^{*}:\;{\boldsymbol{\lambda}}(H_{\boldsymbol{\alpha}})\in\mathbf{R}_{\geq 0}\hbox{ forall }{\boldsymbol{\alpha}}\in R^{+}\}$
be the dominant Weyl chamber in $\mathfrak{h}_{R}^{*}$. For any ${\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C}$, define
$A_{\boldsymbol{\lambda}}:=\sum_{\sigma\in W}\mathrm{sgn}(\sigma)e(\sigma({\boldsymbol{\lambda}})),$
where for any $\sigma\in W$, $\mathrm{sgn}(\sigma)$ is defined to be determinant of $\sigma$ considered as a linear operator on $\mathfrak{h}_{\mathbf{R}}^{*}$. Its value is $1$ or $-1$ since $\sigma$ is a composite of reflections. Let
${\boldsymbol{\rho}}=\frac{1}{2}\sum_{{\boldsymbol{\alpha}}\in R^{+}}{\boldsymbol{\alpha}}$
be half of the sum of positive roots. Then $\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}$ is a trigonometric polynomial function on $T$ invariant under the action of $W$ ([BtD85, VI 1.6]). It defines a continuous function on $T/W\cong G^{\natural}_{\mathbf{R}}$ which we also denote by $\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}$.
The basis $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}$ of $\Gamma$ in Definition 1.3 is also a basis of $\mathfrak{h}_{\mathbf{R}}$, which we fix henceforth. Let $\{\mathbf{e}_{1}^{*},\ldots,\mathbf{e}_{n}^{*}\}$ be the dual basis for $\mathfrak{h}^{*}_{\mathbf{R}}$. For any $\boldsymbol{\lambda}\in\mathfrak{h}^{*}_{\mathbf{R}}$, write $\boldsymbol{\boldsymbol{\lambda}}=\sum_{j=1}^{n}{\lambda}_{j}\mathbf{e}^{*}_{j}$ and let
(1.4.1) ${\mathbf{N}}(\boldsymbol{\lambda}):=(|\lambda_{1}|+1)\cdots(|\lambda_{n}|+1).$
Finally, fix a norm
(1.4.2) $\|\cdot\|:\mathfrak{h}_{\mathbf{R}}\to\mathbf{R}_{\geq 0}$
on $\mathfrak{h}_{\mathbf{R}}.$ The following is a generation of the Erdős–Turán inequality ([Mo94, Chapter 1, Corollary 1.1]) to compact Lie groups.
Theorem 1.5.
For any sequence $x_{1},\ldots,x_{N}\in G_{\mathbf{R}}^{\natural}$, any small box $D$ in $G_{\mathbf{R}}^{\natural}$ and any positive integer $M$, we have
|$\frac{|\{1\leq i\leq N:x_{i}\in D\}|}{N}-\mu_{G_{\mathbf{R}}^{\natural}}(D)$
$\ll$ $\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}c(\boldsymbol{\lambda})\Big{|}\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})\Big{|}+\frac{1}{M},$
where $c(\boldsymbol{\lambda})=\sum_{\sigma\in W}\frac{1}{{\mathbf{N}}(\sigma(\boldsymbol{\lambda}))}$. For any real number $r$, we have
(1.5.1) $\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}c(\boldsymbol{\lambda})\|\boldsymbol{\lambda}\|^{r}\ll M^{r}(\log M)^{n-1}.$
The constants implied by $\ll$ depend only on $G$.
Remark 1.6.
For $G=\mathrm{SL}_{2}$, Theorem 1.5 is due to Niederreiter [Ni91, Lemma 3]. Rouse and Thorner [RT17, Lemma 3.1] give another proof of Niederreiter’s theorem using the Beurling and Selberg majorizing and minorizing functions. We prove Theorem 1.5 using the general theory of the majorizing and minorizing functions of Colzani-Gigante-Travaglini ([CGT11]). In [R13], Rosengarten obtains a version of the Erdős–Turán inequality for simply connected compact Lie groups which is different from ours.
Let $\mathcal{F}^{\prime}$ be another lisse $\overline{\mathbf{Q}}_{\ell}$-sheaves on $X$ punctually $\iota$-pure of weight 00, $\rho_{\mathcal{F}^{\prime}}:\pi_{1}(X,\bar{\eta})\to\mathrm{GL}(\mathcal{F}^{\prime}_{\bar{\eta}})$ the corresponding representation, $G^{\prime}$ the geometric monodromy group, $G^{\prime}_{\mathbf{R}}$ a maximal compact subgroup of $G^{\prime}_{\mathbf{C}}$, $\mu_{G^{\prime\natural}_{\mathbf{R}}}$ the measure on $G^{\prime\natural}_{\mathbf{R}}$ induced by the Haar measure on $G^{\prime}_{\mathbf{R}}$. We have the following theorem on joint distribution.
Theorem 1.7.
Let $\mathcal{F}$ and $\mathcal{F}^{\prime}$ be two lisse $\overline{\mathbf{Q}}_{\ell}$-sheaves punctually $\iota$-pure of weight 00. Suppose their geometric mondromy groups $G$ and $G^{\prime}$ are connected and coincide with their arithmetic monodromy groups, respectively. Suppose furthermore that for any nontrivial irreducible $\overline{\mathbf{Q}}_{\ell}$-representations $\Gamma$ of $G$ and $\Gamma^{\prime}$ of $G^{\prime}$, the representation $(\Gamma\circ\rho_{\mathcal{F}})\otimes(\Gamma^{\prime}\circ\rho_{\mathcal{F}^{\prime}})$ has no nonzero $\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})$-invariant. Then for any small boxes $D$ in $G^{\natural}_{\mathbf{R}}$ and $D^{\prime}$ in $G^{\prime\natural}_{\mathbf{R}}$, we have
|$\ \ \ \ \frac{|\{x\in X(\mathbf{F}_{q^{m}}):\,\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D,\,\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{s}s}\in D^{\prime}\}|}{|X(\mathbf{F}_{q^{m}})|}$
$=\mu_{G^{\natural}_{\mathbf{R}}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})+O\Big{(}q^{-\frac{m}{2(|R^{+}|+|R^{\prime+}|+1)}}(\log q^{m})^{\frac{n+n^{\prime}-2}{|R^{+}|+|R^{\prime+}|+1}}\Big{)},$
where $n$ and $n^{\prime}$ are the dimensions of the Cartan subalgebras of $\mathfrak{g}$ and $\mathfrak{g}^{\prime}$, and the constant implied by $O$ depends only on $X$, $\mathcal{F}$ and $\mathcal{F}^{\prime}$.
As an application, we prove a joint distribution theorem for the Kloosterman sums and the Airy sums. For any $x\in\mathbb{G}_{m}(\mathbf{F}_{q^{m}})\cong\mathbf{F}^{*}_{q^{m}}$, the Kloosterman sum is defined by
$\mathrm{Kl}(\mathbf{F}_{q^{m}},x)=\sum_{z\in\mathbf{F}^{*}_{q^{m}}}e^{\frac{2\pi i}{p}\mathrm{Tr}_{\mathbf{F}_{q^{m}}/\mathbf{F}_{p}}(z+\frac{x}{z})}.$
Deligne [De77, Sommes trig. 7.8] constructs the Kloosterman sheaf $\mathrm{Kl}$ which is a lisse $\overline{\mathbf{Q}}_{\ell}$-adic sheaf on $\mathbb{G}_{m}$ punctually $\iota$-pure of weight $1$ such that
$\iota\mathrm{Tr}(F_{x},\mathrm{Kl}_{\bar{x}})=-\mathrm{Kl}(\mathbf{F}_{q^{m}},x).$
for all $x\in\mathbb{G}_{m}(\mathbf{F}_{q^{m}})$. Consider the twist $\mathrm{Kl}(1/2)$ of $\mathrm{Kl}$ by $\mathbf{Q}_{\ell}(1/2)$ so that $\mathrm{Kl}(1/2)$ is punctually $\iota$-pure of weight 00. Katz ([Ka88, 11.1, 11.3]) shows that both the arithmetic and the geometric monodromy groups of $\mathrm{Kl}(1/2)$ are $\mathrm{SL}_{2}$. Define an angle $\theta(x)$ by
$\mathrm{Kl}(\mathbf{F}_{q^{m}},x)$ $=2q^{m/2}\cos\theta(x).$
For any $x\in\mathbb{A}^{1}(\mathbf{F}_{q^{m}})\cong\mathbf{F}_{q^{m}}$, the Airy sum is defined by
$\mathrm{Ai}(\mathbf{F}_{q^{m}},x)=\sum_{z\in\mathbf{F}_{q^{m}}}e^{\frac{2\pi i}{p}\mathrm{Tr}_{\mathbf{F}_{q^{m}}/\mathbf{F}_{p}}(z^{3}+xz)}.$
Katz ([Ka87, Theorem 17]) constructs the Airy sheaf $\mathrm{Ai}$ which is a lisse $\overline{\mathbf{Q}}_{\ell}$-adic sheaf on $\mathbb{A}^{1}$ punctually pure of weight $1$ such that
$\iota\mathrm{Tr}(F_{x},\mathrm{Ai}_{\bar{x}})=-\mathrm{Ai}(\mathbf{F}_{q^{m}},x).$
for all $x\in\mathbb{A}^{1}(\mathbf{F}_{q^{m}})$. Katz ([Ka87, Theorem 17 (iii), Corollary 20]) proves that $\mathrm{det}(\mathrm{Ai})$ is geometrically constant, and is given by the character $\mathrm{Gal}(\overline{\mathbf{F}}_{q}/\mathbf{F}_{q})\to\overline{\mathbf{Q}}_{\ell}^{*}$ mapping the geometric Frobenius element in $\mathrm{Gal}(\overline{\mathbf{F}}_{q}/\mathbf{F}_{q})$ to $\prod_{\chi^{3}=1,\,\chi\not=1}(-g(\chi,\psi))$, where the product is taken over all nontrivial multiplicative characters $\chi:\mathbf{F}^{*}_{q}\to\overline{\mathbb{Q}}_{\ell}^{*}$ of order dividing $3$ and
$g(\chi,\psi)=\sum_{x\in\mathbf{F}^{*}_{q}}\chi(x)\psi(x)$
is the Gauss sum. Let $\alpha$ be a square root of $\prod_{\chi^{3}=1,\,\chi\not=1}(-g(\chi,\psi))^{-1}$. Let $\mathrm{Ai}(\alpha)$ be the twist of the Airy sheaf by the character $\mathrm{Gal}(\overline{\mathbf{F}}_{q}/\mathbf{F}_{q})\to\overline{\mathbf{Q}}_{\ell}^{*}$ sending the geometric Frobenius element in $\mathrm{Gal}(\overline{\mathbf{F}}_{q}/\mathbf{F}_{q})$ to $\alpha$. Then $\mathrm{Ai}(\alpha)$ is punctually $\iota$-pure of weight 00 and $\mathrm{det}(\mathrm{Ai}(\alpha))=1$. Katz ([Ka87, Theorem 19]) shows that both the arithmetic and the geometric monodromy group of $\mathrm{Ai}(\alpha)$ is $\mathrm{SL}_{2}$ if $p>7$. Define an angle $\theta^{\prime}(x)$ by
$\mathrm{Ai}(\mathbf{F}_{q^{m}},x)$ $=2\alpha^{-m}\cos\theta^{\prime}(x).$
Corollary 1.8.
Suppose $p>7$. For any subintervals $I,I^{\prime}\subset[0,\pi],$ we have
$\frac{1}{q^{m}-1}\Big{|}\{x\in\mathbb{G}_{m}(\mathbf{F}_{q^{m}}):\theta(x)\in I,\;\theta^{\prime}(x)\in I^{\prime}\}\Big{|}=\mu_{\mathrm{ST}}(I)\mu_{\mathrm{ST}}(I^{\prime})+O(q^{-\frac{m}{6}}),$
where $\mu_{ST}=\frac{2}{\pi}\sin^{2}\theta\,\mathrm{d}\theta$ is the Sato-Tate measure on $[0,\pi]$.
Proof.
We apply Theorem 1.7 to the case where $\mathcal{F}=\mathrm{Kl}(1/2)$ and $\mathcal{F}^{\prime}=\mathrm{Ai}(\alpha)$. For $\mathfrak{sl}_{2}$, the dimension of a Cartan subalgebra is
$n=n^{\prime}=1,$
and the number of positive roots is
$|R^{+}|=|R^{\prime+}|=1.$
A maximal compact subgroup of $\mathrm{SL}_{2}(\mathbf{C})$ is $\mathrm{SU}(2)$, and any element in $\mathrm{SU}(2)$ is conjugate to $(\begin{smallmatrix}e^{i\theta}&\\ &e^{-i\theta}\end{smallmatrix})$ for a unique $\theta\in[0,\pi]$. So $\mathrm{SU}(2)^{\natural}$ is identified with $[0,\pi]$. For any $x\in\mathbb{G}_{m}(\mathbf{F}_{q^{m}})$, the conjugacy class $\rho_{\mathrm{Kl}(1/2)}(F_{x})^{\mathrm{ss}}\in\mathrm{SU}(2)^{\natural}$ corresponds to the angle $\theta(x)\in[0,\pi]$, and $\rho_{\mathrm{Ai}(\alpha)}(F_{x})^{\mathrm{ss}}\in\mathrm{SU}(2)^{\natural}$ corresponds to the angle $\theta^{\prime}(x)\in[0,\pi]$. Using Weyl’s integration formula, one can verify $\mu_{\mathrm{SU}(2)^{\natural}}$ is identified with the Sato-Tate measure on $[0,\pi]$. Confer Proposition 3.1 below. Irreducible representations of $\mathrm{SL}_{2}$ are symmetric products of the standard representation. We claim that for any positive integers $k_{1},k_{2}$, $\mathrm{Sym}^{k_{1}}(\mathrm{Kl})\otimes\mathrm{Sym}^{k_{2}}(\mathrm{Ai})$ has no nonzero $\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})$-invariant. Our assertion then follows from Theorem 1.7.
By [Ka88, 4.1.3 and 4.1.4], $\mathrm{Kl}$ is geometrically self dual. So we have
$\ \ \ \ \Big{(}\mathrm{Sym}^{k_{1}}(\mathrm{Kl})_{\bar{\eta}}\otimes\mathrm{Sym}^{k_{2}}(\mathrm{Ai})_{\bar{\eta}}\Big{)}^{\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})}$
$\cong\Big{(}\mathrm{Sym}^{k_{1}}(\mathrm{Kl})_{\bar{\eta}}^{\vee}\otimes\mathrm{Sym}^{k_{2}}(\mathrm{Ai})_{\bar{\eta}}\Big{)}^{\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})}$
$\cong\mathrm{Hom}_{\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})}\Big{(}\mathrm{Sym}^{k_{1}}(\mathrm{Kl})_{\bar{\eta}},\mathrm{Sym}^{k_{2}}(\mathrm{Ai})_{\bar{\eta}}\Big{)}.$
By [De77, Sommes trig. 7.8 (iii)], the sheaf $\mathrm{Sym}^{k_{1}}(\mathrm{Kl})$ is ramified at 00 for all $k_{1}\geq 1$. But $\mathrm{Sym}^{k_{2}}(\mathrm{Ai})$ is unramified at 00. So $\mathrm{Sym}^{k_{1}}(\mathrm{Kl})_{\bar{\eta}}$ and $\mathrm{Sym}^{k_{2}}(\mathrm{Ai})_{\bar{\eta}}$ are non-isomorphic irreducible representations of $\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})$. By Schur’s lemma, we have
$\mathrm{Hom}_{\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})}\Big{(}\mathrm{Sym}^{k_{1}}(\mathrm{Kl})_{\bar{\eta}},\mathrm{Sym}^{k_{2}}(\mathrm{Ai})_{\bar{\eta}}\Big{)}=0.$
So $\mathrm{Sym}^{k_{1}}(\mathrm{Kl})\otimes\mathrm{Sym}^{k_{2}}(\mathrm{Kl})$ has no nonzero $\pi_{1}(\mathbb{G}_{m,\overline{\mathbf{F}}_{q}},\bar{\eta})$-invariant. ∎
The paper is organized as follows. In Section 2, we adopt the work [CGT11] of Colzani, Gigante and Travaglini to our situation. In Section 3, we prove Theorem 1.5 using the Weyl integration formula, the Weyl character formula, and the theorem of Colzani-Gigante-Travaglini. In Section 4 we deduce Theorems 1.2 from Theorem 1.5 and Deligne’s theorem (the Weil conjecture). In Section 4, we study joint distribution and prove Theorem 1.7.
Acknowledgements
We would like to thank Winnie Li and Peng Shan for inspiring discussions. LF is supported by 2021YFA1000700, 2023YFA1009703 and NSFC12171261. YKL is supported by GRF (No. 17303619, 17307720). PX is supported by NSFC (No. 12025106).
2. Trigonometric approximations on the torus
Inspired by Weyl’s criterion of equidistributions and the classical Erdős–Turán inequality on $\mathbf{R},$ Colzani, Gigante and Travaglini [CGT11] obtain a general form of the Erdős–Turán inequality, which covers the torus case ${\mathbf{T}}^{n}:={\mathbf{R}}^{n}/{\mathbf{Z}}^{n}$. For our purpose, we make use of [CGT11, Corollary 1.2 and Corollary 2.9] to derive the following proposition, in which we approximate the characteristic function of a box in $\mathbf{T}^{n}$ by certain trigonometric polynomials.
Proposition 2.1.
Let $I_{j}$ $(j=1,\ldots,n)$ be intervals of lengths $<1$, let $E=\prod_{j=1}^{n}I_{j}$, and let $\Omega$ be the image $E$ in $\mathbf{T}^{n}$. Denote by $\chi_{\Omega}$ the characteristic function of $\Omega.$ For any positive integer $M$, there exist trigonometric polynomials
$B^{\pm}(\mathbf{x})=\sum_{\begin{subarray}{c}\mathbf{k}=(k_{1},\cdots,k_{n})\in\mathbf{Z}^{n}\\ |k_{j}|\leq M\end{subarray}}\widehat{B}^{\pm}(\mathbf{k})e^{2\pi i\mathbf{k}\cdot\mathbf{x}}$
such that
(2.1.1) $B^{-}(\mathbf{x})\leq\chi_{\Omega}(\mathbf{x})\leq B^{+}(\mathbf{x}),$
(2.1.2) $\widehat{B}^{\pm}(\mathbf{0})=\int_{\Omega}\mathrm{d}\mathbf{x}+O\Big{(}\frac{1}{M}\Big{)},\quad\widehat{B}^{\pm}(\mathbf{k})=O\Big{(}\frac{1}{{\mathbf{N}}(\mathbf{k})}\Big{)}\,$
for all nonzero $\mathbf{k}=(k_{1},\cdots,k_{n})\in\mathbf{Z}^{n}$, where
$\mathbf{N}(\mathbf{k}):=\prod_{i=1}^{n}(|k_{i}|+1).$
The constants implied by $O$ in (2.1.2) are independent of $\Omega$.
Proof.
Let $K$ be a nonnegative function on $\mathbf{R}^{n}$ with rapid decay at infinity such that its Fourier transform $\widehat{K}$ satisfies
(2.1.3) $\widehat{K}(\mathbf{0})=1,\quad\widehat{K}(\boldsymbol{\xi})\leq 1\hbox{ if }|\boldsymbol{\xi}|<1,\quad\widehat{K}(\boldsymbol{\xi})=0\hbox{ if }|\boldsymbol{\xi}|\geq 1,$
where $|\boldsymbol{\xi}|=\sqrt{\xi_{1}^{2}+\cdots+\xi_{n}^{2}}$ for any $\boldsymbol{\xi}=(\xi_{1},\cdots,\xi_{n})\in\mathbf{R}^{n}$. Put
$K_{M}(\mathbf{x})$ $:=M^{n}\sum_{\mathbf{k}\in{\mathbf{Z}}^{n}}K(M(\mathbf{x}+\mathbf{k})).$
This defines a periodic function which admits the Fourier expansion
$K_{M}(\mathbf{x})$ $=\sum_{\mathbf{k}\in{\mathbf{Z}}^{n}}\widehat{K}(\mathbf{k}/M)e^{2\pi i\mathbf{k}\cdot\mathbf{x}}.$
According to (2.1.3), $K_{M}$ is a trigonometric polynomial of degree at most $M$ in each variable. By [CGT11, Corollary 1.2] and its proof, one may construct $K$ with the above properties such that
(2.1.4) $K_{M}*\chi_{\Omega}-K_{M}*H_{M}\leq\chi_{\Omega}\leq K_{M}*\chi_{\Omega}+K_{M}*H_{M},$
where
$H_{M}(\mathbf{x})$ $=\frac{1}{4}\psi(2M{\rm dist}(\mathbf{x},\partial E+{\mathbf{Z}}^{n})),$
$\psi(t)$ $=4e^{2\pi}\Big{(}\int_{|\mathbf{x}|\leq 1}K(\mathbf{x})\mathrm{d}\mathbf{x}\Big{)}^{-1}\int_{|\mathbf{x}|\geq t/2}K(\mathbf{x})\,\mathrm{d}\mathbf{x},$
and $0\leq\psi(t)\ll_{\alpha}(1+t)^{-\alpha}$ for any $t\geq 0$ and $\alpha>0$. It follows that the sums
$\sum_{\mathbf{m}\in{\mathbf{Z}}^{n}}\psi(M{\rm dist}(\mathbf{x},\partial E+\mathbf{m}))=\sum_{\mathbf{m}\in{\mathbf{Z}}^{n}}\psi(M{\rm dist}(\mathbf{x}+\mathbf{m},\partial E)),$
are absolutely convergent. We have
$0\leq H_{M}(\mathbf{x})=\frac{1}{4}\psi(2M{\rm dist}(\mathbf{x},\partial E+{\mathbf{Z}}^{n}))\leq\frac{1}{4}\sum_{\mathbf{m}\in{\mathbf{Z}}^{n}}\psi(2M{\rm dist}(\mathbf{x},\partial E+\mathbf{m})).$
Let
$H_{\Omega,M}(\mathbf{x}):=\frac{1}{4}\sum_{\mathbf{m}\in{\mathbf{Z}}^{n}}\psi(2M{\rm dist}(\mathbf{x},\partial E+\mathbf{m})).$
Since $0\leq H_{M}\leq H_{\Omega,M}$ and $K_{M}\geq 0$, we have $0\leq K_{M}*H_{M}\leq K_{M}*H_{\Omega,M}$. Take
$B^{\pm}:=K_{M}*\chi_{\Omega}\pm K_{M}*H_{\Omega,M}.$
The inequalities (2.1.1) then follow from (2.1.4).
The trigonometric polynomials $B^{\pm}$ are explicitly given by
(2.1.5) $B^{\pm}(\mathbf{x})=\sum_{k\in\mathbf{Z}^{n},\,|\mathbf{k}|\leq M}\widehat{K}(\mathbf{k}/M)(\widehat{\chi}_{\Omega}(\mathbf{k})\pm\widehat{H}_{\Omega,M}(\mathbf{k}))e^{2\pi i\mathbf{k}\cdot\mathbf{x}}.$
where $\widehat{K}$ satisfies (2.1.3), $\widehat{\chi}_{\Omega}(\mathbf{k})$ and $\widehat{H}_{\Omega,M}(\mathbf{k})$ are the Fourier coefficients
[…]
$\widehat{H}_{\Omega,M}(\mathbf{k})$ $=\int_{\mathbf{T}^{n}}\frac{1}{4}\sum_{\mathbf{m}\in{\mathbf{Z}}^{n}}\psi(2M{\rm dist}(\mathbf{x},\partial E+\mathbf{m}))e^{-2\pi i\mathbf{k}\cdot\mathbf{x}}\,d\mathbf{x}$
[…]
Note that
(2.1.6) $\widehat{\chi}_{\Omega}(\mathbf{k})$ $=\int_{\Omega}e^{-2\pi i\mathbf{k}\cdot\mathbf{x}}\mathrm{d}\mathbf{x}=\int_{E}e^{-2\pi i\sum_{j=1}^{n}k_{j}x_{j}}\mathrm{d}x_{1}\cdots\mathrm{d}x_{n}$
$=\prod_{j=1}^{n}\int_{I_{j}}e^{-2\pi ik_{j}x_{j}}\mathrm{d}x_{j}\ll\frac{1}{\mathbf{N}(\mathbf{k})}.$
It follows that
$\widehat{K}(\mathbf{0})\widehat{\chi}_{\Omega}(\mathbf{0})=\int_{\Omega}\mathrm{d}\mathbf{x},\quad\widehat{K}(\mathbf{k}/M)\widehat{\chi}_{\Omega}(\mathbf{k})=O\Big{(}\frac{1}{\mathbf{N}(\mathbf{k})}\Big{)}.$
To prove the estimate (2.1.2), it suffices to show
$\widehat{H}_{\Omega,M}(\mathbf{0})=O\Big{(}\frac{1}{M}\Big{)},\quad\widehat{H}_{\Omega,M}(\mathbf{k})=O\Big{(}\frac{1}{\mathbf{N}(\mathbf{k})}\Big{)}$
for any nonzero $\mathbf{k}\in\mathbf{Z}^{n}$ with $|\mathbf{k}|\leq M$, and the implied constants are independent of $\Omega$. By [CGT11, Lemma 2.9], we have
[…]
$\leq$ $c\sum_{j=0}^{n-1}\sum_{A(j)\supset\cdots\supset A(1)}(2M)^{j-n}\prod_{l=1}^{j}\min\bigg{\{}\sqrt{n},\big{(}2\pi\big{|}P_{A(l)}\mathbf{k}\big{|}\big{)}^{-1}\bigg{\}},$
where $c$ is a constant independent of $\Omega$, each $A(l)$ is an $l$ dimensional subspace which is an intersection of a number of codimension 1 subspaces parallel to the faces of the box $E$, $P_{A(l)}\mathbf{k}$ is the orthogonal projection of $\mathbf{k}$ to $A(l)$, and for $j=0$, the inner sum equals the product of $(2M)^{-n}$ and the number of vertices of $E$. Hence the inner sum for $j=0$ is
$(2M)^{-n}2^{n}=M^{-n}.$
For $11\leq j\leq n-$ 1, we have $A(l)={\rm Span}\{\mathbf{e}_{i_{1}},\cdots,\mathbf{e}_{i_{l}}\}$ for some $1\leq i_{1}<\cdots<i_{l}\leq n$, where $\{\mathbf{e}_{1},\cdots,\mathbf{e}_{n}\}$ is the standard basis. Clearly we have
$|P_{A(l)}\mathbf{k}|\geq|k_{i}|\hbox{ for all }i\in\{i_{1},\ldots,i_{\ell}\}.$
Hence for each chain $A(j)={\rm Span}\{e_{i_{1}},\cdots,e_{i_{j}}\}\supset\cdots\supset A(1)$, we have
$(2M)^{j-n}\prod_{l=1}^{j}\min\bigg{\{}\sqrt{n},\big{(}2\pi\big{|}P_{A(l)}\mathbf{k}\big{|}\big{)}^{-1}\bigg{\}}$ $\leq(2M)^{j-n}\prod_{l=1}^{j}\big{(}\sqrt{n}(|k_{i_{l}}|+1)^{-1}\big{)}$
$\leq n^{j/2}\mathbf{N}(\mathbf{k})^{-1},$
where for the last inequality we use the fact that $|k_{i}|+1\leq 2M$ for all $i$. For $\mathbf{k}\neq 0$, we have
$\sum_{j=0}^{n-1}\sum_{A(j)\supset\cdots\supset A(1)}(2M)^{j-n}\prod_{l=1}^{j}\min\bigg{\{}\sqrt{n},\big{(}2\pi\big{|}P_{A(l)}\mathbf{k}\big{|}\big{)}^{-1}\bigg{\}}$
$\leq$ $M^{-n}+\sum_{j=1}^{n-1}\frac{n!}{(n-j)!}n^{j/2}\mathbf{N}(\mathbf{k})^{-1}\ll\mathbf{N}(\mathbf{k})^{-1}.$
For $\mathbf{k}=0$, we have
$(2M)^{j-n}\prod_{l=1}^{j}\min\bigg{\{}\sqrt{n},(2\pi\big{|}P_{A(l)}0\big{|})^{-1}\bigg{\}}=(2M)^{j-n}n^{j/2}\leq n^{j/2}(2M)^{-1}$
since $j\leq n-1$, which yields
$\sum_{j=0}^{n-1}\sum_{A(j)\supset\cdots\supset A(1)}(2M)^{j-n}\prod_{l=1}^{j}\min\bigg{\{}\sqrt{n},\big{(}2\pi\big{|}P_{A(l)}\mathbf{k}\big{|}\big{)}^{-1}\bigg{\}}\ll\frac{1}{M}$
as desired. ∎
3. Proof of Theorem 1.5
Let $\mu_{{G_{\mathbf{R}}}}$ and $\mu_{T}$ be the Haar measures on ${G_{\mathbf{R}}}$ and $T$, respectively. Recall that $\mu_{G_{\mathbf{R}}^{\natural}}$ is the measure on $G_{\mathbf{R}}^{\natural}$ induced by $\mu_{G_{\mathbf{R}}}$. Let $\mu_{T/W}$ be the measure on $T/W$ induced by $\mu_{T}$. Elements in the Weyl group $W$ permute $\{e({\boldsymbol{\alpha}})\}_{{\boldsymbol{\alpha}}\in R}$. So $\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))$ defines a function on $T/W$.
Proposition 3.1.
Via the homeomorphism $\kappa:T/W\stackrel{{\scriptstyle\cong}}{{\to}}G_{\mathbf{R}}^{\natural}$, the measure $\mu_{G_{\mathbf{R}}^{\natural}}$ is identified with the measure $\big{(}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\big{)}\mu_{T/W}$ on $T/W$.
Proof.
Let $\pi:G_{\mathbf{R}}\to G_{\mathbf{R}}^{\natural}$ be the canonical map. For any integrable function $f$ on $G_{\mathbf{R}}^{\natural}$, by the Weyl integration formula [BtD85, IV 1.11, VI 1.8], we have
$\int_{G_{\mathbf{R}}^{\natural}}f\mathrm{d}\mu_{G_{\mathbf{R}}^{\natural}}$ $=\int_{G_{\mathbf{R}}}f\circ\pi\,\mathrm{d}\mu_{G_{\mathbf{R}}}=\frac{1}{|W|}\int_{T}(f\circ\pi|_{T})\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}$
$=\frac{1}{|W|}\int_{T/W}(f\circ\kappa)\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T/W}.$
This completes the proof. ∎
Proposition 3.2.
The set $\{\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}:{\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C}\}$ is an orthonormal basis for the Hilbert space $L^{2}(T/W,(\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}})))\mu_{T/W}).$
Proof.
By [BtD85, Theorem VI 1.7 (ii)], any irreducible representation of $G_{\mathbf{R}}$ has its highest weight lying in $I^{*}\cap\mathfrak{C}$, and for any ${\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C}$, there exists an irreducible representation $\Gamma_{\boldsymbol{\lambda}}$ of $G_{\mathbf{R}}$ unique up to isomorphism with the highest weight ${\boldsymbol{\lambda}}$. By the Peter-Weyl theorem ([BtD85, Theorem III 3.1]), the characters of $\Gamma_{\boldsymbol{\lambda}}$ $({\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C})$ form an orthonormal basis of $L^{2}(G_{\mathbf{R}}^{\natural},\mu_{G_{\mathbf{R}}^{\natural}})$. Identify $(G_{\mathbf{R}}^{\natural},\mu_{G_{\mathbf{R}}^{\natural}})$ with $(T/W,(\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}})))\mu_{T/W})$. By the Weyl character formula [BtD85, VI 1.7 (ii)-(iii)], the function on $T/W$ defined by the character of $\Gamma_{\boldsymbol{\lambda}}$ is exactly $\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}.$
[…]
Lemma 3.3.
For any $\boldsymbol{\lambda}=\sum_{j=1}^{n}\lambda_{j}\mathbf{e}^{*}_{j},$ there exists a constant $c>0$ such that for any positive integer $L$ and any $\sigma\in W$, if $\max_{j}|\lambda_{j}|\leq L$, then $\|\sigma(\boldsymbol{\lambda})\|\leq cL$.
Proof.
It suffices to take $c=\max_{i,\sigma}\sum_{j=1}^{n}\|\sigma(e^{*}_{i})\|$. ∎
For any $\boldsymbol{\lambda}=\sum_{j=1}^{n}\lambda_{j}\mathbf{e}^{*}_{j}\in\mathfrak{h}^{*}_{\mathbf{R}}$, recall that
${\mathbf{N}}(\boldsymbol{\lambda})=(|\lambda_{1}|+1)\cdots(|\lambda_{n}|+1)$
as defined in (1.4.1).
Lemma 3.4.
Let $\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime}\in\mathfrak{h}^{*}_{\mathbf{R}}$. We have
${\mathbf{N}}(\boldsymbol{\lambda}+\boldsymbol{\lambda}^{\prime})\leq{\mathbf{N}}(\boldsymbol{\lambda}){\mathbf{N}}(\boldsymbol{\lambda}^{\prime}),\quad{\mathbf{N}}(\boldsymbol{\lambda}+\boldsymbol{\lambda}^{\prime})^{-1}\ll_{\boldsymbol{\lambda}^{\prime}}{\mathbf{N}}(\lambda)^{-1}.$
Proof.
We have
${\mathbf{N}}(\boldsymbol{\lambda}+\boldsymbol{\lambda}^{\prime})=\prod_{i=1}^{n}(1+|\lambda_{i}+\lambda_{i}^{\prime}|)\leq\prod_{i=1}^{n}\big{(}1+|\lambda_{i}|+|\lambda_{i}^{\prime}|\big{)}\leq{\mathbf{N}}(\boldsymbol{\lambda}){\mathbf{N}}(\boldsymbol{\lambda}^{\prime}).$
Replacing the pair $(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})$ by $(\boldsymbol{\lambda}+\boldsymbol{\lambda}^{\prime},-\boldsymbol{\lambda}^{\prime})$, we get the second assertion. ∎
Lemma 3.5.
Let $I_{j}$ $(j=1,\ldots,n)$ be intervals of lengths $<1$, and let $\Omega$ be the image of the set $\{\sum_{j=1}^{n}t_{j}\mathbf{e}_{j}:\,t_{j}\in I_{j}\}$ under the composite
$\mathfrak{h}_{\mathbf{R}}\to\mathfrak{h}_{\mathbf{R}}/\sum_{j=1}^{n}\mathbf{Z}\mathbf{e}_{j}\underset{\cong}{\stackrel{{\scriptstyle\exp(2\pi i\hbox{-})}}{{\to}}}T.$
We have
$\int_{\Omega}e(\boldsymbol{\lambda})\mathrm{d}\mu_{T}\ll\frac{1}{{\mathbf{N}}(\boldsymbol{\lambda})},$
where the constant implied by $\ll$ depends only on $n$.
Proof.
Via the basis $\{e_{1},\ldots,e_{n}\}$ of $\Gamma$, we may identify $\mathbf{T}^{n}=\mathbf{R}^{n}/\mathbf{Z}^{n}$ with $\mathfrak{h}_{\mathbf{R}}/\Gamma$ and with $T$. Our assertion follows from (2.1.6). ∎
Lemma 3.6.
Let $D$ be a small box in $T/W$, and let $M$ be a positive integer. There exists two linear combinations
$S^{\pm}=\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}\widetilde{S}^{\pm}({\boldsymbol{\lambda}})\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}$
with $\widetilde{S}^{\pm}({\boldsymbol{\lambda}})$ being complex numbers such that the following conditions hold:
  1. (i)
    $S^{\pm}$ are real valued functions on $T/W$, and
    $S^{-}\leq\chi_{D}\leq S^{+}$
    everywhere on $T/W$, where $\chi_{D}$ is the characteristic function of $D$.
  2. (ii)
    $\widetilde{S}^{\pm}(\mathbf{0})=\int_{D}\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T/W}+O\Big{(}\frac{1}{M}\Big{)}.$
  3. (iii)
    For any nonzero ${\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C}$ with $\|\boldsymbol{\lambda}\|\leq M,$ we have
    $\widetilde{S}^{\pm}({\boldsymbol{\lambda}})\ll\sum_{\sigma\in W}\frac{1}{{\mathbf{N}}(\sigma({\boldsymbol{\lambda}}))}.$
Moreover, the implied constants in (ii)-(iii) are independent of $D$.
Proof.
The lattice of integral forms $I^{*}$ can be identified with the group of characters of $T$. Let $R(G_{\mathbf{R}})$ be the Grothendieck ring of representations of $G_{\mathbf{R}}$. The homomorphism
$\mathrm{char}:R(G_{\mathbf{R}})\to\mathbf{Z}[I^{*}]$
sending each representation $\rho$ of $G_{\mathbf{R}}$ to the character of $\rho|_{T}$ induces an isomorphism
$R(G_{\mathbf{R}})\cong\mathbf{Z}[I^{*}]^{W}.$
For any $L^{2}$ function $f$ on $T$, we have the Fourier expansion
$f=\sum_{\boldsymbol{\lambda}\in I^{*}}\widehat{f}(\boldsymbol{\lambda})e(\boldsymbol{\lambda}),$
where
$\widehat{f}(\boldsymbol{\lambda})=\int_{T}fe(-\boldsymbol{\lambda})\mathrm{d}\mu_{T}.$
Choose intervals $I_{j}$ $(j=1,\ldots,n)$ of lengths $<1$ so that $D\subset T/W$ is the image of the set
$\Omega=\Big{\{}\exp\Big{(}2\pi i\sum_{j=1}^{n}t_{j}\mathbf{e}_{j}\Big{)}:\,t_{j}\in I_{j}\Big{\}}$
under the projection $T\to T/W$, and $\Omega\cap\sigma(\Omega)=\emptyset$ for every nonidentity element $\sigma$ in $W$. Let $L=[M/c]$, where $c$ is the constant in Lemma 3.3.
By Theorem 2.1, there exist two trigonometric polynomials
$B^{\pm}=\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\\ \max_{j}|\lambda_{j}|\leq L\end{subarray}}\widehat{B}^{\pm}(\boldsymbol{\lambda})e(\boldsymbol{\lambda})$
such that
(3.6.1) $B^{-}\leq\chi_{\Omega}\leq B^{+},$
(3.6.2) $\widehat{B}^{\pm}(\boldsymbol{0})=\widehat{\chi}_{\Omega}(\boldsymbol{0})+O\big{(}L^{-1}\big{)},\quad\widehat{B}^{\pm}(\boldsymbol{\lambda})=O\Big{(}\frac{1}{{\mathbf{N}}(\boldsymbol{\lambda})}\Big{)},$
for all nonzero $\boldsymbol{\lambda}\in I^{*}$, where $\widehat{\chi}_{\Omega}(\mathbf{0})=\int_{\Omega}\mathrm{d}\mu_{T}$ and the constants implied by $O$ are independent of $\Omega$. Set
$S^{\pm}=\sum_{\sigma\in W}\sigma(B^{\pm}).$
Then $S^{\pm}$ are functions on $T$ invariant under $W$. Hence they define continuous functions on $T/W$ which we denote also by $S^{\pm}$. The inverse image of $D$ under the projection $T\to T/W$ is $\bigsqcup_{\sigma\in W}\sigma(\Omega)$. By (3.6.1), we have
$S^{-}\leq\chi_{D}\leq S^{+}$
as functions on $T/W$. We have $S^{\pm}\in\mathbf{R}[I^{*}]^{W}$. So we can expand $S^{\pm}$ as a finite linear combination of the orthonormal basis $\{\frac{A_{{\boldsymbol{\lambda}}+\boldsymbol{\rho}}}{A_{\boldsymbol{\rho}}}:\,{\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C}\}$ of $L^{2}(T/W,(\frac{1}{|W|}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}})))\mu_{T/W})$. We can write
(3.6.3) $S^{\pm}=\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}\widetilde{S}^{\pm}({\boldsymbol{\lambda}})\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}},$
where
$\widetilde{S}^{\pm}({\boldsymbol{\lambda}})=\frac{1}{|W|}\int_{T}S^{\pm}\overline{\Big{(}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\Big{)}}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}.$
In the expansion (3.6.3), the summation goes over those $\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}$ with $\|\boldsymbol{\lambda}\|\leq M$. This follows from Lemma 3.3, and the fact that in the expansion
$S^{\pm}=\sum_{\sigma\in W}\sigma(B^{\pm})=\sum_{\sigma\in W}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\\ \max_{j}|\lambda_{j}|\leq L\end{subarray}}\widehat{B}^{\pm}(\boldsymbol{\lambda})e(\sigma(\boldsymbol{\lambda})),$
the inner summation goes over those $\boldsymbol{\lambda}\in I^{*}$ with $\max_{j}|\lambda_{j}|\leq L$.
By [BtD85, 1.7 (iii)], we have
$A_{\boldsymbol{\rho}}=e({\boldsymbol{\rho}})\prod_{{\boldsymbol{\alpha}}\in R^{+}}(1-e(-{\boldsymbol{\alpha}})).$
So we have
$A_{\boldsymbol{\rho}}\overline{A_{\boldsymbol{\rho}}}=\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}})),\quad\overline{\Big{(}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\Big{)}}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))=\overline{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}A_{\boldsymbol{\rho}}.$
Using the fact that $\overline{(\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}})}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}$ is $W$-invariant on $T$, we calculate $\widetilde{S}^{\pm}({\boldsymbol{\lambda}})$ as follows:
(3.6.4) $\widetilde{S}^{\pm}({\boldsymbol{\lambda}})$ $=\frac{1}{|W|}\int_{T}S^{\pm}\overline{\Big{(}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\Big{)}}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}$
$=\frac{1}{|W|}\sum_{\sigma\in W}\int_{T}\sigma(B^{\pm})\overline{\Big{(}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\Big{)}}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}$
$=\int_{T}B^{\pm}\overline{\Big{(}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}\Big{)}}\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T}$
$=\int_{T}B^{\pm}\overline{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}A_{\boldsymbol{\rho}}\,\mathrm{d}\mu_{T}$
$=\sum_{\sigma,\tau\in W}\mathrm{sgn}(\sigma\tau)\int_{T}B^{\pm}e(\tau({\boldsymbol{\rho}})-\sigma(\boldsymbol{\rho})-\sigma({\boldsymbol{\lambda}}))\,\mathrm{d}\mu_{T}$
$=\sum_{\sigma,\tau\in W}\mathrm{sgn}(\sigma\tau)\cdot\widehat{B}^{\pm}\big{(}\sigma({\boldsymbol{\lambda}})+\sigma(\boldsymbol{\rho})-\tau({\boldsymbol{\rho}})\big{)}.$
For any nonzero ${\boldsymbol{\lambda}}\in I^{*}\cap\mathfrak{C}$, we claim that $\sigma({\boldsymbol{\lambda}})+\sigma(\boldsymbol{\rho})-\tau({\boldsymbol{\rho}})\not=0$. Otherwise, we have $\tau^{-1}\sigma({\boldsymbol{\lambda}}+\boldsymbol{\rho})=\boldsymbol{\rho}$. Since both $\boldsymbol{\lambda}+\boldsymbol{\rho}$ and $\boldsymbol{\rho}$ lie in the dominant Weyl chamber, we must have $\tau^{-1}\sigma=1$. But then $\boldsymbol{\lambda}+\boldsymbol{\rho}=\boldsymbol{\rho}$ which contradicts to $\boldsymbol{\lambda}\not=0$. So we have
$\widetilde{S}^{\pm}({\boldsymbol{\lambda}})$ $=\sum_{\sigma,\tau\in W}\mathrm{sgn}(\sigma\tau)\cdot\widehat{B}^{\pm}\big{(}\sigma({\boldsymbol{\lambda}})+\sigma(\boldsymbol{\rho})-\tau({\boldsymbol{\rho}})\big{)}$
$\ll\sum_{\sigma,\tau\in W}\frac{1}{{\mathbf{N}}\big{(}\sigma({\boldsymbol{\lambda}})+\sigma(\boldsymbol{\rho})-\tau({\boldsymbol{\rho}})\big{)}}$
$\ll_{\rho,W}\sum_{\sigma}\frac{1}{{\mathbf{N}}(\sigma({\boldsymbol{\lambda}}))},$
where the first inequality follows from (3.6.2), and the second inequality follows from Lemma 3.4. This proves (iii). For ${\boldsymbol{\lambda}}=0$, the calculation (3.6.4) shows that
$\widetilde{S}^{\pm}(\mathbf{0})=\int_{T}B^{\pm}\overline{A_{\boldsymbol{\rho}}}A_{\boldsymbol{\rho}}\,\mathrm{d}\mu_{T}.$
We have
$\Big{|}\widetilde{S}^{\pm}(\mathbf{0})-\int_{D}\frac{1}{|W|}$ $\prod_{{\boldsymbol{\alpha}}\in R}(1-e({\boldsymbol{\alpha}}))\,\mathrm{d}\mu_{T/W}\Big{|}$
$=$ $\Big{|}\int_{T}B^{\pm}\overline{A_{\boldsymbol{\rho}}}A_{\boldsymbol{\rho}}\,\mathrm{d}\mu_{T}-\int_{\Omega}\overline{A_{\boldsymbol{\rho}}}A_{\boldsymbol{\rho}}\,\mathrm{d}\mu_{T}\Big{|}$
$=$ $\Big{|}\int_{T}(B^{\pm}-\chi_{\Omega})\overline{A_{\boldsymbol{\rho}}}A_{\boldsymbol{\rho}}\,\mathrm{d}\mu_{T}\Big{|}$
$\leq$ $\|\overline{A_{\boldsymbol{\rho}}}A_{\boldsymbol{\rho}}\|_{\infty}\int_{T}|B^{\pm}-\chi_{\Omega}|\,\mathrm{d}\mu_{T}.$
By (3.6.2), the last expression is
$\|\overline{A_{\boldsymbol{\rho}}}A_{\boldsymbol{\rho}}\|_{\infty}\int_{T}(B^{+}-\chi_{\Omega})\mathrm{d}\mu_{T}=\|\overline{A_{\boldsymbol{\rho}}}A_{\boldsymbol{\rho}}\|_{\infty}(\widehat{B}^{+}(\mathbf{0})-\widehat{\chi}_{\Omega}(\mathbf{0}))=O\Big{(}\frac{1}{M}\Big{)}$
in the ``+"``"``+"` ` + " case, and
$\|\overline{A_{\boldsymbol{\rho}}}A_{\boldsymbol{\rho}}\|_{\infty}\int_{T}(\chi_{\Omega}-B^{-})\mathrm{d}\mu_{T}=\|\overline{A_{\boldsymbol{\rho}}}A_{\boldsymbol{\rho}}\|_{\infty}(\widehat{\chi}_{\Omega}(\mathbf{0})-\widehat{B}^{-}(\mathbf{0}))=O\Big{(}\frac{1}{M}\Big{)}$
in the ``"``"``-"` ` - " case. ∎
3.7. Proof of Theorem 1.5
By Lemma 3.6, we have
|$\ \ \ \ \frac{|\{1\leq i\leq N:x_{i}\in D\}|}{N}-\mu_{G_{\mathbf{R}}^{\natural}}(D)=\frac{1}{N}\sum_{i=1}^{N}\chi_{D}(x_{i})-\mu_{G_{\mathbf{R}}^{\natural}}(D)$
$\leq\frac{1}{N}\sum_{i=1}^{N}S^{+}(x_{i})-\mu_{G_{\mathbf{R}}^{\natural}}(D)$
$=\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}\widetilde{S}^{+}({\boldsymbol{\lambda}})\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})+\widetilde{S}^{+}(\mathbf{0})-\mu_{G_{\mathbf{R}}^{\natural}}(D)$
$\ll\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}c(\boldsymbol{\lambda})\Big{|}\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})\Big{|}+\frac{1}{M},$
where
$c(\boldsymbol{\lambda})=\sum_{\sigma\in W}\frac{1}{{\mathbf{N}}(\sigma({\boldsymbol{\lambda}}))}.$
This gives the desired upper bound. We get the lower bound using $S^{-}$.
It remains to prove (1.5.1). For any $\lambda=\sum_{j=1}^{n}\lambda_{j}\mathbf{e}_{j}^{*}$, let $\|\boldsymbol{\lambda}\|_{\infty}=\max_{j}|\lambda_{j}|$. Taking $N:=M\sup\{\|\sigma(\boldsymbol{\lambda})\|_{\infty}:\,\sigma\in W,\;\|\boldsymbol{\lambda}\|=1\}$, we have
$\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}c(\boldsymbol{\lambda})\|\boldsymbol{\lambda}\|^{r}$ $=\sum_{\sigma\in W}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}\frac{\|\boldsymbol{\lambda}\|^{r}}{{\mathbf{N}}(\sigma({\boldsymbol{\lambda}}))}$
$\leq\sum_{\sigma\in W}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}-\{\mathbf{0}\}\\ \max_{j}|\lambda_{j}|\leq N\end{subarray}}\frac{\|\sigma(\boldsymbol{\lambda})\|^{r}}{{\mathbf{N}}({\boldsymbol{\lambda}})}.$
Note that
$\|\sigma(\boldsymbol{\lambda})\|=\Big{\|}\sum_{j}\lambda_{j}\sigma(\mathbf{e}_{j}^{*})\Big{\|}\ll\max_{j}|\lambda_{j}|.$
We thus have
$\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}-\{\mathbf{0}\}\\ \max_{j}|\lambda_{j}|\leq N\end{subarray}}\frac{\|\sigma(\boldsymbol{\lambda})\|^{r}}{{\mathbf{N}}({\boldsymbol{\lambda}})}$ $\ll\sum_{\begin{subarray}{c}\lambda_{1},\ldots,\lambda_{n}\in\mathbf{Z}\\ |\lambda_{1}|\leq\cdots\leq|\lambda_{n}|\leq N\end{subarray}}\frac{(|\lambda_{n}|+1)^{r-1}}{(|\lambda_{1}|+1)\cdots(|\lambda_{n-1}|+1)}$
$\ll N^{r}(\log N)^{n-1}\ll M^{r}(\log M)^{n-1}.$
This completes the proof.
4. Proof of Theorem 1.2
Throughout this section, $X$ is a smooth geometrically connected scheme over $\mathbf{F}_{q}$, and $\eta$ is the generic point of $X$. Any lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf $\mathcal{G}$ on $X$ corresponds to a representation $\rho_{\mathcal{G}}:\pi_{1}(X,\bar{\eta})\to\mathrm{GL}(\mathcal{G}_{\bar{\eta}})$. We say $\mathcal{G}$ has no geometric invariant (resp. no geometric coinvariant) if
$\mathcal{G}_{\bar{\eta}}^{\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})}=0\quad(\hbox{resp. }\mathcal{G}_{\bar{\eta},\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})}=0).$
Choose a finite extension $E$ of $\mathbf{Q}_{\ell}$ such that $\mathcal{G}$ comes from lisse a $E$-sheaf $\mathcal{G}_{E}$. Let $R$ be the discrete valuation ring of $E$, $\kappa$ the residue field, and $\mathcal{G}_{R}$ an $R$-sheaf such that $\mathcal{G}_{E}=\mathcal{G}_{R}\otimes_{R}E$. Choose a finite Galois étale covering $X^{\prime}\to X$ so that $(\mathcal{G}_{R}\otimes_{R}\kappa)|_{X^{\prime}}$ is a constant sheaf.
Proposition 4.1.
Let $\mathcal{G}$ be a lisse $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$ punctually $\iota$-pure of weight 00, and let $d=\mathrm{dim}\,X$. Suppose $\mathcal{G}$ has no geometric invariant. We have
$\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\iota\mathrm{Tr}(F_{x},\mathcal{G}_{\bar{x}})\Big{|}$ $\leq C\,q^{\frac{(2d-1)m}{2}}\mathrm{dim}(\mathcal{G}_{\bar{\eta}}),$
where $C$ is a constant depending only on the Galois étale covering $X^{\prime}\to X$ described above.
Proof.
Let $\mathcal{G}^{\vee}=\mathcal{H}om(\mathcal{G},\overline{\mathbf{Q}}_{\ell})$ be the dual sheaf of $\mathcal{G}$. By the Poincaré duality theorem, we have
$H^{2d}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})\cong\mathrm{Hom}\Big{(}H^{0}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G}^{\vee}),\overline{\mathbf{Q}}_{\ell}(-d)\Big{)}$
$\cong$ $\mathrm{Hom}\Big{(}\mathcal{G}_{\bar{\eta}}^{\vee,\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})},\overline{\mathbf{Q}}_{\ell}(-d)\Big{)}\cong\mathcal{G}_{\bar{\eta},\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})}(-d).$
Since $\mathcal{G}$ is punctually $\iota$-pure of weight 00, it is geometrically semisimple by [De80, 3.4.1(iii)]. Since $\mathcal{G}$ has no geometric invariant, it has no geometric coinvariant and hence
$H^{2d}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})=0.$
Combined with the Grothendieck trace formula [De77, Rapport 3.2], we get
$\sum_{x\in X(\mathbf{F}_{q^{m}})}\mathrm{Tr}(F_{x},\mathcal{G}_{\bar{x}})=\sum_{i=0}^{2d-1}\mathrm{Tr}\Big{(}F^{m},H^{i}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})\Big{)},$
where $F$ is the geometric Frobenius correspondence. By Deligne’s theorem [De80, 3.3.1], all eigenvalues of $F$ on $H^{i}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})$ have absolute values $\leq q^{\frac{i}{2}}$ via $\iota$. By [FW04, Lemma 1.8], there exists a constant $c$ depending only on the Galois étale covering $X^{\prime}\to X$ such that
$\mathrm{dim}\,H^{i}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})\leq c\,\mathrm{dim}(\mathcal{G}_{\bar{\eta}}).$
So we have
$\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\iota\mathrm{Tr}(F_{x},\mathcal{G}_{\bar{x}})\Big{|}$$\leq$ $\sum_{i=0}^{2d-1}q^{\frac{i}{2}}\,\mathrm{dim}\,H^{i}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})$
$\leq$ $2cdq^{\frac{(2d-1)m}{2}}\mathrm{dim}(\mathcal{G}_{\bar{\eta}}).$
Remark 4.2.
Suppose $X$ is a smooth geometrically connected algebraic curve. By [De80, 1.4.1], we have
$H^{0}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})$$\cong\begin{cases}\mathcal{G}_{\bar{\eta}}^{\pi_{1}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\bar{\eta})}&\text{if }X\text{ is proper},\\ 0&\text{otherwise}.\end{cases}$
As $\mathcal{G}$ has no geometric invariant, we have $H^{0}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})=0$ and hence $H^{1}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})$ is the only cohomology group which may not vanish. By the Grothendieck–Ogg–Shafarevich formula [Gr77, X 7.12], we have
$\mathrm{dim}\,H^{1}_{c}(X\otimes_{\mathbf{F}_{q}}\overline{\mathbf{F}}_{q},\mathcal{G})$ $=(2g-2+N)\mathrm{dim}(\mathcal{G}_{\bar{\eta}})+\sum_{x\in\overline{X}-X}\mathrm{deg}(x)\mathrm{sw}_{x}(\mathcal{G}).$
So we have
$\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\iota\mathrm{Tr}(F_{x},\mathcal{G}_{\bar{x}})\Big{|}$ $\leq q^{\frac{m}{2}}\Big{(}(2g-2+N)\mathrm{dim}(\mathcal{G}_{\bar{\eta}})+\sum_{x\in\overline{X}-X}\mathrm{deg}(x)\mathrm{sw}_{x}(\mathcal{G})\Big{)}.$
This is more explicit than the estimate in Proposition 4.1.
4.3. Proof of Theorem 1.2
For any nonzero $\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}$, let
$\Gamma_{\boldsymbol{\lambda}}:G\to\mathrm{GL}(r,\overline{\mathbf{Q}}_{\ell})$
be the irreducible representation of $G$ with the highest weight $\boldsymbol{\lambda}$, where $r=\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})$. Let $\mathcal{G}_{\boldsymbol{\lambda}}$ be the $\overline{\mathbf{Q}}_{\ell}$-sheaf corresponding to the representation
$\pi_{1}(X,\bar{\eta})\stackrel{{\scriptstyle\rho_{\mathcal{F}}}}{{\to}}G\stackrel{{\scriptstyle\Gamma_{\boldsymbol{\lambda}}}}{{\to}}\mathrm{GL}(r,\overline{\mathbf{Q}}_{\ell}).$
Since $\Gamma_{\boldsymbol{\lambda}}$ is nontrivial and irreducible, the sheaf $\mathcal{G}_{\boldsymbol{\lambda}}$ has no geometric invariant. Moreover, $\mathcal{G}_{\boldsymbol{\lambda}}$ is punctually $\iota$-pure of weight 00 since $\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}$ lies in the maximal compact subgroup $G_{\mathbf{R}}$ of $G$ for every closed point $x$ in $X$. By the Weyl character formula, we have
$\iota\mathrm{Tr}(F_{x},\mathcal{G}_{\boldsymbol{\lambda},\bar{x}})=\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}).$
We claim that there is a common finite extension $E$ of $\mathbf{Q}_{\ell}$ and a finite Galois étale covering $X^{\prime}\to X$ such that for each $\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}$, there exists an $R$-sheaf $\mathcal{G}_{\boldsymbol{\lambda},R}$ such that $\mathcal{G}_{\boldsymbol{\lambda}}$ comes from the $E$-sheaf $\mathcal{G}_{\boldsymbol{\lambda},R}\otimes_{R}E$ and $\mathcal{G}_{\boldsymbol{\lambda},R}\otimes_{R}\kappa|_{X^{\prime}}$ is a constant sheaf. Indeed, we can first choose such a common pair $(E,X^{\prime}\to X)$ for a finite family of weights $\boldsymbol{\lambda}_{1},\ldots,\boldsymbol{\lambda}_{m}$ which form a family of generators for the monoid $I^{*}\cap\mathfrak{C}$. Then the pair $(E,X^{\prime}\to X)$ works for all $\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}$ since $\Gamma_{\boldsymbol{\lambda}}$ can be realized as a sub-representation of a representation constructed by taking tensor products of representations from the family $\Gamma_{\boldsymbol{\lambda}_{1}},\ldots,\Gamma_{\boldsymbol{\lambda}_{m}}$. Applying Proposition 4.1 to the sheaf $\mathcal{G}_{\boldsymbol{\lambda}}$, we get
(4.3.1) $\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}})\Big{|}$ $=\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\iota\mathrm{Tr}(F_{x},\mathcal{G}_{\boldsymbol{\lambda},\bar{x}})\Big{|}$
$\ll\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})q^{\frac{(2d-1)m}{2}},$
By the Weyl dimension formula [BtD85, VI 1.7 (iv)], we have
$\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})=\prod_{{\boldsymbol{\alpha}}\in R^{+}}\frac{({\boldsymbol{\lambda}}+{\boldsymbol{\rho}})(H_{\boldsymbol{\alpha}})}{{\boldsymbol{\rho}}(H_{\boldsymbol{\alpha}})}.$
It follows that
(4.3.2) $\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})\ll\|\boldsymbol{\lambda}\|^{|R^{+}|}$
for any nonzero $\boldsymbol{\lambda}$. Combining (4.3.1) and (4.3.2) with Theorem 1.5, for any positive integer $M$, we have
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D\}|}{|X(\mathbf{F}_{q^{m}})|}-\mu_{G_{\mathbf{R}}^{\natural}}(D)$
$\ll$ $\frac{1}{|X(\mathbf{F}_{q^{m}})|}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}c(\boldsymbol{\lambda})\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}})\Big{|}+\frac{1}{M}$
$\ll$ $\frac{1}{q^{dm}}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C}-\{\mathbf{0}\}\\ \|\boldsymbol{\lambda}\|\leq M\end{subarray}}c(\boldsymbol{\lambda})\|\boldsymbol{\lambda}\|^{|R^{+}|}q^{\frac{(2d-1)m}{2}}+\frac{1}{M}$
$\ll$ $q^{-\frac{m}{2}}M^{|R^{+}|}(\log M)^{n-1}+\frac{1}{M}.$
Taking $M$ to be the integer closest to $q^{\frac{m}{2(|R^{+}|+1)}}(\log q^{m})^{-\frac{n-1}{|R^{+}|+1}}$. Then $\log M\ll\log q^{m}$, and we have
$q^{-\frac{m}{2}}M^{|R^{+}|}(\log M)^{n-1}+\frac{1}{M}$
$\ll$ $q^{-\frac{m}{2}}M^{|R^{+}|}(\log q^{m})^{n-1}+\frac{1}{M}$
$\ll$ $q^{-\frac{m}{2}}q^{\frac{m{|R^{+}|}}{2(|R^{+}|+1)}}(\log q^{m})^{-\frac{(n-1){|R^{+}|}}{|R^{+}|+1}}(\log q^{m})^{n-1}+q^{-\frac{m}{2(|R^{+}|+1)}}(\log q^{m})^{\frac{n-1}{|R^{+}|+1}}$
$=$ $2q^{-\frac{m}{2(|R^{+}|+1)}}(\log q^{m})^{\frac{n-1}{|R^{+}|+1}}$
So we have
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D\}|}{|X(\mathbf{F}_{q^{m}})|}=\mu_{G_{\mathbf{R}}^{\natural}}(D)+O\Big{(}q^{-\frac{m}{2(|R^{+}|+1)}}(\log q^{m})^{\frac{n-1}{|R^{+}|+1}}\Big{)}.$
5. Joint distribution
Denote data associated to $G^{\prime}$ by symbols with a superscript . For example, for each $\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}$, let $\lambda^{\prime}_{j}$ be the coefficients in the expression of $\boldsymbol{\lambda}^{\prime}$ as a linear combination of the basis $\{\mathbf{e}_{1}^{*},\ldots,\mathbf{e}^{*}_{n^{\prime}}\}$ of $\mathfrak{h}^{\prime*}_{\mathbf{R}}$ dual to the basis $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}$ of $\mathfrak{h}^{\prime}_{\mathbf{R}}$ in the definition of small boxes for $G^{\prime\natural}$. Let $\|\cdot\|^{\prime}$ be the chosen norm on $\mathfrak{h}^{\prime*}_{\mathbf{R}}$.
Proposition 5.1.
For any sequences $x_{1},\ldots,x_{N}\in G_{\mathbf{R}}^{\natural}$ and $x^{\prime}_{1},\ldots,x^{\prime}_{N}\in G^{\prime\natural}_{\mathbf{R}}$, any small boxes $D$ in $G_{\mathbf{R}}^{\natural}$ and $D^{\prime}$ in $G^{\prime\natural}_{\mathbf{R}}$, and any positive integer $M$, we have
|$\frac{|\{1\leq i\leq N:x_{i}\in D,\,x^{\prime}_{i}\in D^{\prime}\}|}{N}-\mu_{G_{\mathbf{R}}^{\natural}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})$
$\ll\frac{1}{N}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(\mathbf{0},\mathbf{0})\\ \|\boldsymbol{\lambda}\|,\|\boldsymbol{\lambda}^{\prime}\|^{\prime}\leq M\end{subarray}}c(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\Big{|}\sum_{i=1}^{N}\frac{A_{{\boldsymbol{\lambda}}+{\boldsymbol{\rho}}}}{A_{\boldsymbol{\rho}}}(x_{i})\frac{A_{{\boldsymbol{\lambda}^{\prime}}+{\boldsymbol{\rho}^{\prime}}}}{A_{\boldsymbol{\rho}^{\prime}}}(x^{\prime}_{i})\Big{|}+\frac{1}{M},$
where
$c(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})=\sum_{\sigma\in W,\,\sigma^{\prime}\in W^{\prime}}\frac{1}{{\mathbf{N}}(\lambda){\mathbf{N}}(\lambda^{\prime})}.$
For any real numbers $r$ and $r^{\prime}$, we have
(5.1.1) $\quad\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(\mathbf{0},\mathbf{0})\\ \|\boldsymbol{\lambda}\|,\|\boldsymbol{\lambda}^{\prime}\|^{\prime}\leq M\end{subarray}}c(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})(\|\boldsymbol{\lambda}\|+1)^{r}(\|\boldsymbol{\lambda}^{\prime}\|^{\prime}+1)^{r^{\prime}}\ll M^{r+r^{\prime}}(\log M)^{n+n^{\prime}-2}$
The implied constants depend only on $G$ and $G^{\prime}$.
Proof.
The first part is just Theorem 1.5 for the group $G\times G^{\prime}$. We only need to verify (5.1.1) which is stronger than applying Theorem 1.5 directly by saving a logarithm factor. This is due to the special feature of the action of the Weyl group $W\times W^{\prime}$.
Following the arguments in the proof of proof of Theorem 1.5, we have
$\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(\mathbf{0},\mathbf{0})\\ \|\boldsymbol{\lambda}\|,\|\boldsymbol{\lambda}^{\prime}\|^{\prime}\leq M\end{subarray}}c(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})(\|\boldsymbol{\lambda}\|+1)^{r}(\|\boldsymbol{\lambda}^{\prime}\|^{\prime}+1)^{r^{\prime}}$
=$\sum_{\begin{subarray}{c}\sigma\in W\\ \sigma^{\prime}\in W^{\prime}\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(\mathbf{0},\mathbf{0})\\ \|\boldsymbol{\lambda}\|,\|\boldsymbol{\lambda}^{\prime}\|^{\prime}\leq M\end{subarray}}\frac{(\|\boldsymbol{\lambda}\|+1)^{r}(\|\boldsymbol{\lambda}^{\prime}\|^{\prime}+1)^{r^{\prime}}}{{\mathbf{N}}(\sigma({\boldsymbol{\lambda}})){\mathbf{N}}(\sigma^{\prime}({\boldsymbol{\lambda}^{\prime}}))}$
$\ll\sum_{\begin{subarray}{c}\sigma\in W\\ \sigma^{\prime}\in W^{\prime}\end{subarray}}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(\mathbf{0},\mathbf{0})\\ \max_{j,j^{\prime}}\{|\lambda|_{j},|\lambda^{\prime}|_{j^{\prime}}\}\leq N\end{subarray}}\frac{(\|\sigma(\boldsymbol{\lambda})\|+1)^{r}(\|\sigma^{\prime}(\boldsymbol{\lambda}^{\prime})\|+1)^{r^{\prime}}}{{\mathbf{N}}({\boldsymbol{\lambda}}){\mathbf{N}}({\boldsymbol{\lambda}^{\prime}})},$
where $N=M\sup\{\|\sigma(\boldsymbol{\lambda})\|_{\infty},\|\sigma^{\prime}(\boldsymbol{\lambda}^{\prime})\|_{\infty}:\sigma\in W,\sigma^{\prime}\in W^{\prime},\|\boldsymbol{\lambda}\|=\|\boldsymbol{\lambda}^{\prime}\|^{\prime}=1\}$. (Confer the proof Theorem 1.5 for the notation.) We have
$\|\sigma(\boldsymbol{\lambda})\|+1\ll\max_{1\leq j\leq n}|\lambda_{j}|+1,\quad\|\sigma^{\prime}(\boldsymbol{\lambda}^{\prime})\|^{\prime}+1\ll\max_{1\leq j\leq n}|\lambda^{\prime}_{j}|+1.$
We infer that
$\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(\mathbf{0},\mathbf{0})\\ \max_{j,j^{\prime}}\{|\lambda|_{j},|\lambda^{\prime}|_{j^{\prime}}\}\leq N\end{subarray}}\frac{(\|\sigma(\boldsymbol{\lambda})\|+1)^{r}(\|\sigma^{\prime}(\boldsymbol{\lambda}^{\prime})\|^{\prime}+1)^{r^{\prime}}}{{\mathbf{N}}({\boldsymbol{\lambda}}){\mathbf{N}}({\boldsymbol{\lambda}^{\prime}})}$
$\ll\sum_{\begin{subarray}{c}\lambda_{1},\ldots,\lambda_{n},\lambda^{\prime}_{1},\ldots,\lambda^{\prime}_{n^{\prime}}\in\mathbf{Z}\\ |\lambda_{1}|\leq\cdots\leq|\lambda_{n}|\leq N\\ |\lambda^{\prime}_{1}|\leq\cdots\leq|\lambda^{\prime}_{n^{\prime}}|\leq N\end{subarray}}\frac{(|\lambda_{n}|+1)^{r-1}}{(|\lambda_{1}|+1)\cdots(|\lambda_{n-1}|+1)}\frac{(|\lambda^{\prime}_{n^{\prime}}|+1)^{r^{\prime}-1}}{(|\lambda’_{1}|+1)\cdots(|\lambda^{\prime}_{n^{\prime}-1}|+1)}$
$\ll N^{r+r^{\prime}}(\log N)^{n+n^{\prime}-2}\ll M^{r+r^{\prime}}(\log M)^{n+n^{\prime}-2}.$
This completes the proof. ∎
5.2. Proof of Theorem 1.7
For any $\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}$, let $\Gamma_{\boldsymbol{\lambda}^{\prime}}$ be the irreducible representation of $G^{\prime}$ with the highest weight $\boldsymbol{\lambda}^{\prime}$, and let $\mathcal{G}_{\boldsymbol{\lambda}^{\prime}}$ be the $\overline{\mathbf{Q}}_{\ell}$-sheaf on $X$ corresponding to the representation $\Gamma_{\boldsymbol{\lambda}^{\prime}}\circ\rho_{\mathcal{F}^{\prime}}$. By our assumption, if $(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(0,0)$, the sheaf $\mathcal{G}_{\boldsymbol{\lambda}}\otimes\mathcal{G}_{\boldsymbol{\lambda}^{\prime}}$ has no geometric invariant and is punctually $\iota$-pure of weight 00. By the Weyl character formula, we have
$\iota\mathrm{Tr}\Big{(}F_{x},(\mathcal{G}_{\boldsymbol{\lambda}}\otimes\mathcal{G}_{\boldsymbol{\lambda}^{\prime}})_{\bar{x}}\Big{)}=\frac{A_{\boldsymbol{\lambda}+\boldsymbol{\rho}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}})\frac{A_{\boldsymbol{\lambda}^{\prime}+\boldsymbol{\rho}^{\prime}}}{A_{\boldsymbol{\rho}^{\prime}}}(\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{ss}}).$
Applying Proposition 4.1 to the sheaf $\mathcal{G}_{\boldsymbol{\lambda}}\otimes\mathcal{G}_{\boldsymbol{\lambda}^{\prime}}$, we get
$\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\frac{A_{\boldsymbol{\lambda}+\boldsymbol{\rho}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}})\frac{A_{\boldsymbol{\lambda}^{\prime}+\boldsymbol{\rho}^{\prime}}}{A_{\boldsymbol{\rho}^{\prime}}}(\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{ss}})\Big{|}\ll\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}^{\prime}})q^{\frac{(2d-1)m}{2}}.$
Note that
$\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}})\mathrm{dim}(\Gamma_{\boldsymbol{\lambda}^{\prime}})\ll(\|\boldsymbol{\lambda}\|+1)^{|R^{+}|}(\|\boldsymbol{\lambda}^{\prime}\|^{\prime}+1)^{|R^{\prime+}|}.$
By Proposition 5.1, for any positive integer $M$, we have
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\,\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D,\,\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{s}s}\in D^{\prime}\}|}{|X(\mathbf{F}_{q^{m}})|}-\mu_{G_{\mathbf{R}}^{\natural}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})$
$\ll$ $\frac{1}{|X(\mathbf{F}_{q^{m}})|}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(\mathbf{0},\mathbf{0})\\ \|\boldsymbol{\lambda}\|,\|\boldsymbol{\lambda}^{\prime}\|^{\prime}\leq M\end{subarray}}c(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\Big{|}\sum_{x\in X(\mathbf{F}_{q^{m}})}\frac{A_{\boldsymbol{\lambda}+\boldsymbol{\rho}}}{A_{\boldsymbol{\rho}}}(\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}})\frac{A_{\boldsymbol{\lambda}^{\prime}+\boldsymbol{\rho}^{\prime}}}{A_{\boldsymbol{\rho}^{\prime}}}(\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{ss}})\Big{|}+\frac{1}{M}$
$\ll$ $\frac{1}{q^{dm}}\sum_{\begin{subarray}{c}\boldsymbol{\lambda}\in I^{*}\cap\mathfrak{C},\,\boldsymbol{\lambda}^{\prime}\in I^{\prime*}\cap\mathfrak{C}^{\prime}\\ (\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})\not=(\mathbf{0},\mathbf{0})\\ \|\boldsymbol{\lambda}\|,\|\boldsymbol{\lambda}^{\prime}\|^{\prime}\leq M\end{subarray}}c(\boldsymbol{\lambda},\boldsymbol{\lambda}^{\prime})(\|\boldsymbol{\lambda}\|+1)^{|R^{+}|}(\|\boldsymbol{\lambda}^{\prime}\|^{\prime}+1)^{|R^{\prime+}|}q^{\frac{(2d-1)m}{2}}+\frac{1}{M}$
$\ll$ $q^{-\frac{m}{2}}M^{|R^{+}|+|R^{\prime+}|}(\log M)^{n+n^{\prime}-2}+\frac{1}{M}.$
Taking $M$ to be the integer closest to $q^{\frac{m}{2(|R^{+}|+|R^{\prime+}|+1)}}(\log q^{m})^{-\frac{n+n^{\prime}-2}{|R^{+}|+|R^{\prime+}|+1}}$. Then $\log M\ll\log q^{m}$, and we have
$q^{-\frac{m}{2}}M^{|R^{+}|+|R^{\prime+}|}(\log M)^{n+n^{\prime}-2}+\frac{1}{M}\ll$ $q^{-\frac{m}{2}}M^{|R^{+}|+|R^{\prime+}|}(\log q^{m})^{n+n^{\prime}-2}+\frac{1}{M}$
$\ll$ $q^{-\frac{m}{2(|R^{+}|+|R^{\prime+}|+1)}}(\log q^{m})^{\frac{n+n^{\prime}-2}{|R^{+}|+|R^{\prime+}|+1}}$
So we have
|$\frac{|\{x\in X(\mathbf{F}_{q^{m}}):\,\rho_{\mathcal{F}}(F_{x})^{\mathrm{ss}}\in D,\,\rho_{\mathcal{F}^{\prime}}(F_{x})^{\mathrm{s}s}\in D^{\prime}\}|}{|X(\mathbf{F}_{q^{m}})|}-\mu_{G_{\mathbf{R}}^{\natural}}(D)\mu_{G^{\prime\natural}_{\mathbf{R}}}(D^{\prime})$
$=$ $O\Big{(}q^{-\frac{m}{2(|R^{+}|+|R^{\prime+}|+1)}}(\log q^{m})^{\frac{n+n^{\prime}-2}{|R^{+}|+|R^{\prime+}|+1}}\Big{)}.$